### v1

# ELDeR: Getting Efficient LLMs through Data-Driven Regularized Layer-wise Pruning

Abstract
The deployment of Large language models (LLMs) in many fields is largely hindered by their high computational and memory costs. Recent studies suggest that LLMs exhibit sparsity, which can be used for pruning. Previous pruning methods typically follow a prune-then-finetune paradigm. Since the pruned parts still contain valuable information, statically removing them without updating the remaining parameters often results in irreversible performance degradation, requiring costly recovery fine-tuning (RFT) to maintain performance. To address this, we propose a novel paradigm: first apply regularization, then prune. Based on this paradigm, we propose ELDeR: Getting Efficient LLMs through Data-Driven Regularized Layer-wise Pruning. We multiply the output of each transformer layer by an initial weight, then we iteratively learn the weights of each transformer layer by using a small amount of data in a simple way. After that, we apply regularization to the difference between the output and input of the layers with smaller weights, forcing the information to be transferred to the remaining layers. Compared with direct pruning, ELDeR reduces the information loss caused by direct parameter removal, thus better preserving the model’s language modeling ability. Experimental results show that ELDeR achieves superior performance compared with powerful layer-wise structured pruning methods, while greatly reducing RFT computational costs. Since ELDeR is a layer-wise pruning method, its end-to-end acceleration effect is obvious, making it a promising technique for efficient LLMs.
1 Introduction
Large language models (LLMs) have made remarkable progress in various tasks and fields, showcasing their strong capabilities [1, 2, 3, 4]. As models grow larger, the increasing number of parameters results in substantial computational and memory demands, posing major challenges to the practical deployment of LLMs. There is an urgent need for techniques that can compress model size without compromising performance.
The main methods for reducing model size include pruning [5, 6, 7], quantization [8, 9], distillation [10], and low-rank decomposition [11]. This study primarily concentrates on pruning, an effective and widely adaptable technique that is orthogonal to other model compression methods. Pruning methods are typically divided into two main types: unstructured pruning [12, 13] and structured pruning [14, 15]. In contrast to unstructured pruning, structured pruning enables recovery fine-tuning (RFT) for specific downstream tasks without the need for specialized hardware. It can also achieve significant end-to-end acceleration without the need for specially designed hardware [16].
Despite advancements, existing structured pruning methods have some limitations. They all follow the paradigm of first prune channels or layers based on a designed metric, and then performing RFT. However, they overlook that important information may also exist in the pruned parts [17, 18, 19], and directly removing them without updating the remaining parameters often results an irreversible decline in model performance. Therefore, obtaining satisfactory performance with pruned models typically requires a large amount of data for RFT [14], increasing computational overhead. Moreover, high sparsity can cause the model’s performance to collapse, limiting the acceleration effect.
To address these limitations, we present a new pruning paradigm that first apply regularization and followed by pruning. The regularization process dynamically transfers crucial information from the layers to be pruned to the remaining layers, thus minimizing the information loss resulting from direct parameter removal. This paradigm can maintain good model performance without RFT. Based on this paradigm, we propose ELDeR, a simple and effective data-driven regularized layer-wise pruning method. The comparison of the ELDeR with existing layer-wise pruning approaches is shown in Figure 1. Specifically, the process of ELDeR is divided into the following steps: (1) select a small amount of data for learning layer weights and regularization, (2) iteratively learn the layer weights, (3) apply regularization to the difference between the outputs and inputs of the layers to be pruned, (4) prune. First, we randomly choose a small subset of data from commonly used benchmark datasets. Since the selected data is small, the computational costs in learning the layer weights and regularization process can be greatly minimized. Second, we iteratively learn the weights of each layer by adding their $\ell_{1}$-norm in the loss function. Third, we apply regularization ($\ell_{1}$-norm or $\ell_{2}$-norm) to the difference between the outputs and inputs of the layers with smaller weights, forcing important information to be transferred to the remaining layers which significantly reduces the performance decline caused by parameter removal. Thus the model can maintain good performance even at high pruning ratios, thereby providing significant end-to-end acceleration. Finally, we prune the layers with smaller weights. Extensive experiments demonstrate that ELDeR substantially outperforms current layer-wise pruning methods in terms of perplexity and accuracy across different pruning ratios, while also significantly improving end-to-end acceleration. The main contributions of ELDeR are summarized as follows:
  • Propose A New Paradigm: ELDeR reduces information loss caused by direct parameter removal by sequentially learning layer weights, applying regularization ($\ell_{1}$-norm or $\ell_{2}$-norm), and performing pruning. This approach helps preserve model performance without requiring for RFT.
  • High Performance: ELDeR outperforms previous layer-wise structured pruning methods in both generation and zero-shot tasks. The pruned model demonstrates a considerable acceleration compared to the dense model.
  • Low Overhead: By using only a small amount of data for learning layer weights and regularization, ELDeR significantly reduces the overhead of RFT.
Figure 1: A comparison between existing layer-wise structured pruning methods and ELDeR. Deeper blue layer represents greater performance impact, the taller cylinder represents larger data volume.
2 Related Works
2.1 Pruning Methods
The process of removing unnecessary weights, known as pruning, has proven to be an effective method for reducing the complexity of deep neural networks over the past few decades [20, 21, 22, 23]. Pruning methods can generally be divided into two categories: unstructured pruning [24, 25, 26] and structured pruning [5, 27, 28].
Unstructured pruning methods eliminate individual weights according to their significance. For instance, Magnitude [22, 29] assumes larger weights are more important, Wanda [13] uses the product of the layer weights and activations as the criterion, and SparseGPT [12] uses the weight squared divided by the inverse Hessian matrix. Their main advantage is flexibility. However, their drawbacks include the need for specialized hardware to accelerate [30], the requirement for extreme sparsity to achieve considerable acceleration [31], and are not suitable for RFT on downstream tasks.
Structured pruning methods, including channel-wise pruning [15] and layer-wise pruning [32]. Channel-wise pruning methods typically create a metric to assess the importance of channels in the parameter matrixs of LLMs, and then remove the less important ones. Layer-wise pruning methods treat entire layers as the basic units for pruning. For instance, SLEB [16], iteratively remove entire transformer layers by evaluating their impact on the model’s final loss, ShortGPT [33] thinks high similarity between layers means redundancy, LaCo [27] uses layer collapse to prune, Shortened LLaMA [34] prune layers in one-shot based on their importance. However, these methods have a major limitation: even less important channels or layers may still contain valuable information, and directly pruning them often results in significant performance degradation. To overcome this, we explore the possibility of "shifting" important information from the pruned parts to the remaining parts, which could greatly maintain model performance.
2.2 Regularization
Regularization is commonly used in machine learning [35, 36, 37], including preventing model overfitting [38] and feature selection [39]. The $\ell_{1}$-norm typically promotes sparsity by driving certain coefficients to zero, while the $\ell_{2}$-norm encourages smoother solutions [40]. Both can significantly affect the distribution pattern of the data [22, 41]. Inspired by this, we propose a new pruning paradigm, in which the regularization process could transfer important information from the pruned parameter space to the remaining parts of the model, thus improving performance.
3 Methodology
In this section, we provide a detailed introduction to ELDeR. The overall ELDeR algorithm is summarized in Algorithm 1. ELDeR mainly consists of four steps:
  • Select Data: Select a small amount of data to learn the layer weights and apply regularization.
  • Learn Layer Weights: Iteratively learn the weights of each layer by adding their $\ell_{1}$-norm to the loss function.
  • Regularization: Apply regularization to the difference between the output and input of layers with smaller weights, facilitating information transfer.
  • Pruning: Removes the layers that were regularized in the previous step.
3.1 Select Data
The data is selected from widely used datasets, including Alpaca [42], WikiText-2 [43], PTB [44], and C4 [45]. For instance, approximately 128 samples are randomly selected from the WikiText-2 training dataset for learning layer weights and applying regularization.
3.2 Learn Layer Weights
For clarity, we first define the relevant notations. $\mathbf{W}$ represents the initial model. Let $p$ be the pruning ratio, indicating that $p\%$ of model parameters will be pruned. The number of layers in the initial model is denoted by $l$. The model hidden size is $d$. Let $\mathbf{X}\in\mathbb{R}^{b\times n\times d}$ represents the input data embedding, where $b$ is the batch size and $n$ is the number of tokens. The input to the $i$th layer is denoted as $\mathbf{X_{\text{in}}^{i}}\in\mathbb{R}^{b\times n\times d}$, and the output from the $i$th layer is denoted as $\mathbf{X_{\text{out}}^{i}}\in\mathbb{R}^{b\times n\times d}$. A learnable weight $S[i]$ is assigned to the $i$th transformer layer, and the set of all layer weights is denoted as $S$.
Algorithm 1 ELDeR algorithm.
1:Input: selected data: $\mathbf{X}$, number of layers in LLM: $l$, initial model $\mathbf{W}$, number of layers to prune: $n$, set of layer weights: $S$, set of layers to be pruned: $P$, norm type: flag.
2:Initialize each layer weight in $S$ to 1.
3:$P\leftarrow\emptyset$, $MinS\leftarrow 1e9$
4:for $i=0$ to $n-1$ do
5:     $S\leftarrow learnWeights(\mathbf{W},S,\mathbf{X})$
6:     for $j=0$ to $l-i-1$ do
7:         if $S[j]<MinS$ then
8:              $MinS\leftarrow S[j]$
9:              $MinS\_id\leftarrow j$
10:         end if
11:     end for
12:     $\mathbf{W}\leftarrow mask(\mathbf{W},MinS\_id)$
13:     $P\leftarrow P\bigcup\{MinS\_id\}$
14:end for
15:$\mathcal{L}_{\text{sum}}\leftarrow\mathcal{L}(\mathbf{W},\mathbf{X})$
16:for $i=0$ to $\mathrm{sizeof}(P)$ do
17:     $\mathcal{L}_{\text{sum}}\leftarrow\mathcal{L}_{\text{sum}}+regularized(\mathbf{W}[P[i]])$
18:end for
19:update $\mathbf{W}$ using backpropagation algorithm
20:$\mathbf{W}\leftarrow Prune(\mathbf{W},P)$
According to Algorithm 1, we initialize the weight of each layer to 1. As shown in Figure 2, the output of each layer is scaled by its associated weight before being passed as input to the subsequent layer. To learn the weight of each layer, we employ the input data embedding $\mathbf{X}$. When pruning $n$ layers, $\mathbf{X}$ is repeatedly used as input in each iteration. The objective function defined in Equation 1 comprises two parts: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$, and the sum of the $\ell_{1}$-norm of all layer weights, $\lambda_{1}$ balances the two components, $l_{1}$ is the number of layers in the current model that are not masked.
$\begin{split}\mathcal{L}_{\text{learn}}=\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{1}\sum_{i=0}^{l_{1}-1}\|S[i]\|_{1}\end{split}$ (1)
Forward and backward propagation is then performed to learn the set of layer weights $S$. Subsequently, the layer with the smallest weight is identified and masked out in the next iteration, and its index is added to the pruning set $P$. This process is iteratively performed $n$ times if there are $n$ layers need to be pruned which follows a greedy strategy. We also explored a one-shot pruning approach, in which a single forward and backward propagation is used to identify the $n$ layers with the lowest weights. However, as shown in Section 4.8, this often results in the removal of consecutive transformer layers, which leads to a substantial degradation in model performance.
We formalize the minimization of the objective function in Equation 1 as an optimization problem. Since $\ell_{1}$-norm is not differentiable, backpropagation (BP) can’t be used directly, we need to transform the problem using Proposition 3.1.
Proposition 3.1.
(If the loss function includes the $\ell_{1}$-norm, it can also be solved using BP. Proof in Appendix A). The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
$\min\quad||x||_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (2)
s.t. $\quad-y\leq x\leq y,$
$\quad y\geq 0.$
The optimization problem in Equation 1 can be equivalently transformed into a constrained formulation in Equation 3. After transformation, the objective function is differentiable, and can be solved by BP.
$\min_{W,y}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{1}\mathbf{1}^{T}y$ (3)
s.t. $\quad-y\leq S\leq y,$
$\quad y\geq 0.$
3.3 Regularization
After completing learning the layer weighs, we further lighten the impact of the pruned layers on the
Figure 2: Details of layer weight learning and regularization. $P$ is the set of prunable layers, identified iteratively using the method in Section 3.2.
final model output by applying regularization to them. The approach is simple and straightforward: we use the input data embedding $\mathbf{X}$ as input and apply a one-shot regularization on the difference between the output and input of each layer in the pruning set $P$. For example, in Figure 2, the $i$th layer in $P$ and should be regularized, we just add the $\ell_{1}$-norm or $\ell_{2}$-norm of $\mathbf{X_{out}^{i}-X_{in}^{i}}$ to the loss function. This encourages the information to be redistributed from the pruned layers to the remaining ones, thereby reducing the amount of information retained in the pruned layers and significantly minimizing the performance degradation caused by subsequent pruning.
The objective function defined in Equation 4 comprises two components: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$ and the regularization loss, $\lambda_{2}$ balances the two components. When the $\ell_{1}$-norm is used, the equivalent constrained optimization problem transformed using Proposition 3.1 is shown in Equation 5. This formulation can be directly solved using BP.
$\begin{split}\mathcal{L}_{\text{sum}}=\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{2}\sum_{i=0}^{|P|-1}\|\mathbf{X_{out}^{i}-X_{in}^{i}}\|\end{split}$ (4)
$\min_{W,Y_{i}}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{2}\sum_{i=0}^{|P|-1}\mathbf{1}^{T}\mathbf{Y_{i}}\mathbf{1}^{T}$ (5)
s.t. $\quad-\mathbf{Y_{i}}\leq\mathbf{X_{\text{out}}^{i}-X_{\text{in}}^{i}}\leq\mathbf{Y_{i}},$
$\quad\mathbf{Y_{i}}\geq 0.$
3.4 Pruning
After applying the regularization, we directly remove the transformer layers in the set $P$, which is a straightforward process.
4 Experiments
This section introduces experimental setup (4.1) and analyzes the effectiveness of ELDeR from the following aspects: performance comparison (4.2), acceleration 4.3, robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (4.7), choice of learning layer weights (4.8), impact of regularization (4.9).
4.1 Experimental Setup
Implementation: All methods are implemented in PyTorch [46], using the Hugging Face Transformers library [47]. All experiments are conducted on 80GB NVIDIA A100 GPUs. For fairness, we use llm-eval-harness [48] to evaluate the pruned models. More details are in Appendix B.
Datasets: We evaluated on generation and zero-shot tasks. For generation task, following prior work[15], we evaluate the model’s perplexity on WikiText-2 [43] test set. For zero-shot task, we evaluate on PIQA [49], WinoGrande [50], HellaSwag [51], ARC-e and ARC-c [52]. In Section 4.5 we use a randomly selected subset of data from Alpaca [42], WikiText-2 [43], PTB [44], and C4 [45].
Models: The models pruned using ELDeR include the LLaMA models (LLaMA2-7B, LLaMA2-13B, LLaMA3-8B) [1, 53], OPT models (OPT-2.7B, OPT-13B) [54], and Phi-2 [55].
Baselines: We evaluate ELDeR against competitive layer-wise structured pruning methods: SLEB [16], ShortGPT [33], LaCo [27], Shortened LLaMA (PPL version) [34]. We also compared with powerful unstructured pruning methods like Magnitude [22], Wanda [13], and SparseGPT [12] in Appendix D.
Evaluation Metrics: The performance on generation task is measured by perplexity, a well-established and robust metric [56], while zero-shot tasks performance is evaluated using accuracy [57]. The acceleration effects are represented by throughput and latency [16].
4.2 Performance Comparison
For fairness, we randomly selected 128 samples from the WikiText-2 training dataset with a sequence length of 2048 as the data for ELDeR to get layer weights and perform regularization, as well as the calibration data for the baselines. Following [15], we selected another 1,000 samples from WikiText-2 training dataset to perform RFT leveraging LoRA [58] on the baselines after pruning. The aforementioned 128 samples and the 1,000 samples have no overlap. Because ELDeR is finetuning-free, we did not perform RFT on ELDeR. The pruning ratio was 25% which means 25% of the transformer layers in a model will be pruned.
Table 1: Performance comparison of different pruning methods. ‘PPL’ refers to perplexity on WikiText-2. The accuracy is reported on five zero-shot benchmarks. The best result is in bold, and the second-best is underlined. ELDeR-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, ELDeR-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm. ‘PR’ is the pruning ratio.
Model Method PR PPL ($\downarrow$) PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 SLEB 25% 7.65 68.85 63.63 50.96 49.38 30.79 52.72 ShortGPT 25% 7.15 69.67 65.19 51.26 52.46 33.89 54.49 LaCo 25% 7.38 68.53 63.76 50.39 51.28 33.45 53.48 Shortened LLaMA 25% 7.74 66.88 62.19 51.45 51.47 32.66 52.93 ELDeR-$\ell_{2}$ 25% 6.53 71.35 67.62 55.84 52.26 35.75 56.56 ELDeR-$\ell_{1}$ 25% 6.58 71.42 67.13 56.05 52.18 35.79 56.51 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 SLEB 25% 15.71 65.19 56.26 44.54 46.28 25.14 47.48 ShortGPT 25% 14.96 67.37 57.65 46.82 49.43 26.54 49.56 LaCo 25% 15.38 66.54 59.45 43.68 48.74 24.96 48.67 Shortened LLaMA 25% 15.89 63.14 57.36 43.57 47.62 25.88 47.51 ELDeR-$\ell_{2}$ 25% 13.18 70.54 60.27 46.35 51.59 27.36 51.22 ELDeR-$\ell_{1}$ 25% 13.12 70.65 60.13 46.24 51.76 27.55 51.27 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 SLEB 25% 9.63 65.22 63.38 55.51 56.39 33.46 54.79 ShortGPT 25% 8.89 66.75 66.26 57.14 58.93 36.42 57.10 LaCo 25% 9.14 69.45 65.31 52.67 55.73 34.89 55.61 Shortened LLaMA 25% 9.47 65.58 64.72 58.36 54.19 32.96 55.16 ELDeR-$\ell_{2}$ 25% 7.08 72.48 67.52 60.45 62.69 39.73 60.57 ELDeR-$\ell_{1}$ 25% 7.17 72.08 67.93 60.42 62.38 39.89 60.54 LLaMA3-8B Dense 0% 5.76 85.56 77.94 79.27 78.84 56.49 75.62 SLEB 25% 10.38 72.74 64.12 67.74 65.84 45.16 63.12 ShortGPT 25% 9.26 75.38 69.25 70.12 68.54 47.57 66.17 LaCo 25% 10.14 74.69 67.52 66.36 69.43 46.31 64.86 Shortened LLaMA 25% 9.84 73.62 68.73 68.55 66.16 43.39 64.09 ELDeR-$\ell_{2}$ 25% 7.84 77.25 71.63 72.26 71.49 49.58 68.44 ELDeR-$\ell_{1}$ 25% 7.68 77.36 71.33 72.82 70.75 49.66 68.38 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 SLEB 25% 11.96 72.83 64.06 63.32 59.98 34.65 58.97 ShortGPT 25% 11.38 73.59 64.52 65.68 60.41 34.99 59.84 LaCo 25% 11.79 73.96 63.24 62.17 61.46 33.28 58.82 Shortened LLaMA 25% 11.62 71.26 63.57 66.29 58.47 33.83 58.68 ELDeR-$\ell_{2}$ 25% 10.45 74.15 64.47 68.82 61.55 35.23 60.84 ELDeR-$\ell_{1}$ 25% 10.32 74.58 64.37 68.45 61.29 34.87 60.71 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 SLEB 25% 7.08 68.31 66.86 57.12 62.19 38.45 58.59 ShortGPT 25% 6.79 73.26 68.37 62.25 67.54 43.38 62.96 LaCo 25% 7.14 72.35 65.78 59.43 65.36 42.73 61.13 Shortened LLaMA 25% 6.92 69.45 66.38 59.86 65.25 39.12 60.01 ELDeR-$\ell_{2}$ 25% 5.82 74.56 69.34 64.79 71.25 45.63 65.11 ELDeR-$\ell_{1}$ 25% 5.89 75.06 69.18 64.96 71.43 45.26 65.18
As shown in Table 1, ELDeR achieves the lowest perplexity and the highest average accuracy across all models compared to the baseline methods, demonstrating its superior performance on both generation and zero-shot tasks. On LLaMA2-7B, its perplexity is 20% lower than the second-best method, ShortGPT. Notably, on OPT-13B, compared to the dense model, ELDeR only drop 1% in average accuracy. This further demonstrates the effectiveness of ELDeR. The performance difference between $\ell_{2}$-norm and $\ell_{1}$-norm is tiny, indicating that ELDeR is insensitive to the choice of regularization norm. In the following sections, we refer to ELDeR as the version that uses the $\ell_{2}$-norm.
4.3 Acceleration
Language processing in LLMs consists of two main stages: prompt processing, which is compute-bound, and token generation, which is memory-bound. We analyze the speedup achieved in each stage separately. Table 2 shows the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated on a single 80GB NVIDIA A100 GPU. Following the methodology of previous work [16], the token generation test involves generating 128-token sentences with a batch size of 64, while prompt processing latency is measured using an input sequence of 2048 tokens.
Table 2: Throughput (tokens/s) and latency (ms) on OPT-13B and LLaMA2-13B. ‘PPL’ refers to the perplexity on Wikitext2. ‘TI’ is Throughput Increase.
Model Method Pruning ratio PPL($\downarrow$) Avg_Acc(%) Tokens/s($\uparrow$) TI($\uparrow$) Latency($\downarrow$) Speedup($\uparrow$) OPT-13B Dense 0% 10.12 61.79 1029 1.00× 386.5 1.00× ELDeR 25% 10.45 60.84 1348 1.31× 286.3 1.35× ELDeR 50% 15.38 50.83 1801 1.75× 208.9 1.85× LLaMA2-13B Dense 0% 4.88 71.76 1066 1.00× 396.9 1.00× ELDeR 25% 5.82 65.11 1386 1.30× 298.4 1.33× ELDeR 50% 11.28 57.57 1823 1.71× 216.9 1.83×
At a 50% pruning ratio on OPT-13B, ELDeR achieves a 75% increase in throughput and a 46% reduction in latency compared to the dense model. For LLaMA2-13B, it delivers a 71% improvement in throughput and a 45% decrease in latency. These results underscore ELDeR’s superior efficiency in accelerating model inference.
4.4 Robustness to Different Pruning Ratios
Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%. The perplexity of LLaMA2-7B under different methods are shown in Figure 4. ELDeR significantly outperforms other methods across various pruning ratios. At a pruning ratio of 20%, ELDeR performs similarly to the dense model. When the pruning ratio is up to 60%, SLEB collapses, while ELDeR maintains relatively low perplexity compared to other pruning methods. This demonstrates that ELDeR maintains robust performance even under extreme pruning ratios, enabling structured pruning to unlock significant potential for model acceleration. More detailed results are in Appendix E.
Figure 3: Perplexity of LLaMA2-7B pruned by five methods under different pruning ratios on WikiText-2.
Figure 4: Comparison of perplexity on different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
4.5 Dependency on Datasets
Since ELDeR relies on data-driven regularization, we investigate its dataset dependency. Keeping all other settings consistent with Section 4.2, we only change the source of the 128 samples: they are drawn respectively from Alpaca, WikiText-2, PTB, and C4. We evaluated perplexity of five methods on WikiText-2. As shown in Figure 4, ELDeR consistently outperforms the other methods across datasets, demonstrating its robustness.
4.6 Low Overhead
We evaluated the perplexity of baselines on LLaMA2-7B using varying amounts of data for RFT, keeping all other conditions consistent with Section 4.2 and only change the data size from 1,000 to 8,000. ELDeR is finetuing-free. From Figure 5, it can be observed that ELDeR outperforms the other methods with approximately 4.000 RFT data, significantly reducing data overhead. We speculate the reason is that ELDeR iteratively learns layer weights and then applies regularization, allowing it to more accurately identify layers to prune before pruning, and subsequently transfer the information from those layers to the remaining parts of the model through regularization. This process reduces information loss caused by pruning, thus preserving model performance and lowering the cost associated with RFT.
4.7 Ablation Study
Effect of Learning Layer Weights Iteratively ELDeR learns layer weights in a greedy and iterative manner. As shown in Table 3 (Row 3 and Row 6), replacing iterative layer weight learning with a one-shot approach leads to increased model perplexity and decreased accuracy, highlighting the importance of learning layer weights iteratively, which will be discussed in detail in Section 4.8.
Effect of Applying Regularization As shown in Table 3 (Row 4 and Row 7), removing the regularization process results in increased model perplexity and decreased accuracy, demonstrating the effectiveness of applying regularization, which will be discussed in detail in Section 4.9.
Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda_{1}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$ and $\lambda_{2}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. We perform a grid search over $\lambda_{1}$ and $\lambda_{2}$, with details provided in Appendix C. The optimal combination yielding the lowest perplexity on LLaMA2-7B is $\lambda_{1}=5\times 10^{-3}$ and $\lambda_{2}=10^{-3}$.
Figure 5: The comparison of perplexity on WikiText-2 between ELDeR and other baseline methods under different amounts of RFT data.
Table 3: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘w/o W’ refers learning layer weights in one-shot, ‘w/o R’ refers no regularization. The pruning ratio is 25%.
Model Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B ELDeR 7.08 0.00 60.57 0.00 w/o W 9.26 +2.18 56.19 -4.38 w/o R 10.15 +3.07 54.36 -6.21 LLaMA2-13B ELDeR 5.82 0.00 65.11 0.00 w/o W 8.35 +2.53 59.36 -5.75 w/o R 9.47 +3.65 56.25 -8.86
4.8 Choice of Learning Layer Weights
In this section, we explore the effectiveness of (1) iteratively learning layer weights using a greedy strategy compared to (2) acquiring all layer weights at one-shot. We set the pruning ratio to 25% and conduct experiments on LLaMA2-7B (32 layers), keeping all settings consistent with Section 4.2 except for the method used to determine layer weights.
The perplexity and average accuracy of the pruned models are shown in Table 4. It can be observed that, compared to using a greedy strategy to iteratively learn layer weights, learning all layer weights in one-shot exhibits significant degradation in model performance. This behavior can be explained by the observation that the importance of a block changes as other blocks are removed. The results of selecting the eight least important layers using both methods are shown in Figure 6. It can be seen that learning layer weights in one-shot tends to select consecutive layers. While these blocks may individually have limited impact on LLM inference performance, removing a continuous sequence of blocks can significantly degrade the overall inference results.
Table 4: The pruned LLaMA2-7B performance under different learning layer weights methods.
Cases PPL($\downarrow$) Avg_Acc(%) (1) 7.08 60.57 (2) 9.26 56.19
Figure 6: Results of selecting the eight lowest-weight layers using iterative and one-shot layer weight learning.
4.9 Impact of Regularization
We keep other settings consistent with Section 4.2. After learning the layer weights iteratively, we consider two scenarios: (1) without regularization and (2) with regularization, followed by pruning. The perplexity and average accuracy under (1) and (2) on LLaMA2-7B are shown in Table 5. The model exhibits lower perplexity and higher average accuracy with regularization. Since perplexity is the exponential form of cross-entropy loss, lower perplexity corresponds to a lower cross-entropy loss. A smaller loss function indicates better model performance indicating that the regularization process mitigates the impact of pruning on the overall model.
According to previous work [59], the output and input of each layer in LLMs exhibit high similarity. This inherent similarity enables us to apply regularization on the difference between the input and output of certain layers using only a small portion of data. We compute the cosine similarity according to Equation 6 on the LLaMA2-7B model between the input representations $\mathbf{X_{\text{in}}^{i}}\in\mathbb{R}^{b\times n\times d}$ and the output representations $\mathbf{X_{\text{out}}^{i}}\in\mathbb{R}^{b\times n\times d}$, where $b$ denotes the batch size, $n$ the number of tokens, and $d$ is the hidden size of the model. The computation is performed on layers [3, 6, 9, 11, 13, 20, 23, 29], which are selected using the iterative method described in Section 4.8, both before and after regularization.
$\text{CosSim}(\mathbf{X_{\text{in}}^{i}},\mathbf{X_{\text{out}}^{i}})=\frac{1}{bn}\sum_{k=1}^{b}\sum_{j=1}^{n}\frac{\left\langle\mathbf{X_{\text{in}}^{i}}[k,j,:],\ \mathbf{X_{\text{out}}^{i}}[k,j,:]\right\rangle}{\left\|\mathbf{X_{\text{in}}^{i}}[k,j,:]\right\|_{2}\cdot\left\|\mathbf{X_{\text{out}}^{i}}[k,j,:]\right\|_{2}}$ (6)
As shown in Figure 7, the input-output similarity of these layers is already high before regularization and increases even further after regularization. The increased similarity in the layers with regularization indicates that the input undergoes less change after passing through these layers, suggesting that less information is retained in them. In the Appendix F, we illustrate the changes in similarity between the input and output of the layers that were not regularized, before and after regularization. The results in Figure 8 of Appendix F show that the similarity in these layers decreases, meaning that, after regularization, the input undergoes greater changes when passing through these layers than before. This demonstrates that the regularization process weakens the influence of the regularized layers and enhances the influence of the unregularized parts, suggesting that regularization process may facilitate the transfer of information from the regularized layers to the rest of the model.
Table 5: The performance differences between (1) and (2).
Cases PPL($\downarrow$) Avg_Acc(%) (1) 7.08 60.57 (2) 10.15 54.36
Figure 7: The input-output similarity of the regularized layers.
5 Conclusion
In this paper, we propose a new pruning paradigm: first apply regularization, then prune. Unlike previous paradigm that first prune and then apply RFT, ELDeR transfers critical information from the pruned parameter space to the remaining parts during regularization, effectively mitigating the irreversible performance degradation caused by information loss. ELDeR demonstrates superior performance in generation and zero-shot tasks, significantly outperforming existing methods. For instance, ELDeR surpasses the powerful ShotGPT by 20% in perplexity on LLaMA2-7B. On OPT-13B, under 25% sparsity, the average accuracy drops by only 1%, while achieving 1.35× speedup compared to the dense model. Moreover, ELDeR is finetuning-free, substantially reducing computational costs and the reliance on RFT. The new paradigm may provide insights for structured pruning in LLMs.
Appendix A Proof of Proposition 3.1
Step 1: Expressing $\ell_{1}$-norm Using Elements.
The objective function in the unconstrained problem is the $\ell_{1}$-norm of the vector $x$, which is defined as:
$||x||_{1}=\sum_{i=1}^{n}|x_{i}|$ (7)
This function aims to minimize the sum of the absolute values of the components of $x$.
Step 2: Reformulating the Constrained Problem
The constrained optimization problem introduces an auxiliary variable $y$, where for each element $i$:
$x_{i}\geq-y_{i}\quad\text{and}\quad x_{i}\leq y_{i}\quad$ (8)
This implies that $y_{i}\geq|x_{i}|$, meaning each element of $y$ serves as an upper bound for the absolute value of the corresponding element in $x$. Consequently, minimizing $|x|_{1}$ is equivalent to minimizing the sum of the elements in $y$. Thus, the objective function is defined as:
$\mathbf{1}^{T}y$ (9)
Thus, minimizing $\mathbf{1}^{T}y$ is equivalent to minimizing the sum of the absolute values of $x$, which is the $\ell_{1}$-norm of $x$.
This transformation allows the optimization problem to be solved without directly involving the absolute value function, resulting in an equivalent constrained optimization problem that can be addressed via backpropagation. Thus, the proof of the Proposition 3.1 is complete.
Appendix B Detailed Implementation
In this part, we first introduce several hyperparameter settings, with the detailed results shown in Table 6. In our experiments, we employ FP16 precision for all evaluated models, including Phi-2, OPT-2.7B, LLaMA3-8B, OPT-13B, LLaMA2-7B, and LLaMA2-13B. For all RFT configurations, we set the LoRA rank $r$ to 32, the scaling factor $\alpha$ to 10, and the sequence length to 2048. All other hyperparameters follow the default settings provided in the Hugging Face PEFT package [60]. We set the batch size to 64. In future work, we will further explore a broader range of batch sizes.
Table 6: Implementation Details
Precision LoRA Rank Scaling Factor Max Sequence Length Batch Size Learning Rate Early Stop Threshold Min Delta FP16 32 10 2048 64 2e-5 5 1e-4
To ensure a fair comparison between ELDeR and other methods, we maintain consistency in the data used across all approaches. Specifically, the data used by ELDeR for learning layer weights and performing regularization is identical to the calibration data used by the baseline methods. Furthermore, we ensure that the data employed during the RFT process is consistent across all baseline methods, thereby enabling a controlled and equitable evaluation framework. Following previous works [16], for the comparison unstructured pruning methods like Mangnitude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by ELDeR for learning weights and regularization.
Appendix C Optimal $\lambda_{1}$ and $\lambda_{2}$ for LLaMA2-7B
Keeping all other settings consistent with Section 4.2, we evaluate the model’s perplexity on WikiText-2 by varying $\lambda_{1}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$ and $\lambda_{2}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$, resulting in a total of 49 combinations.
It can be observed from Table 7 that the optimal combination yielding the lowest perplexity on LLaMA2-7B is $\lambda_{1}=5\times 10^{-3}$ and $\lambda_{2}=10^{-3}$.
When fixing $\lambda_{2}=10^{-3}$, we gradually decrease $\lambda_{1}$ from $5\times 10^{-3}$ to $10^{-5}$, during which the model’s perplexity increases steadily. This suggests that when the $\ell_{1}$-norm loss constitutes a relatively small portion of the total loss during the iterative learning of layer weights, it fails to effectively constrain the layer weights, resulting in larger deviations. Conversely, when $\lambda_{1}$ is gradually increased from $5\times 10^{-3}$ to $10^{-1}$, the model’s perplexity also increases, indicating that a dominant $\ell_{1}$-norm loss in the total objective function can hinder the optimization of the language modeling capability.
As shown in Table 7, when $\lambda_{1}=5\times 10^{-3}$ is fixed, setting $\lambda_{2}=10^{-5}$ results in a regularization loss that is too weak to effectively redistribute important information. This leads to a substantial increase in perplexity after pruning. On the other hand, when $\lambda_{2}=10^{-1}$, the overly strong regularization impairs the model’s language modeling capability, also resulting in a noticeable drop in performance after pruning. The best performance is observed at $\lambda_{2}=10^{-3}$, highlighting the critical need to balance the language modeling loss and regularization loss.
Table 7: The optimal $\lambda_{1}$ and $\lambda_{2}$ for LLaMA2-7B.
$\lambda_{1}$ / $\lambda_{2}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ $5\times 10^{-2}$ $10^{-1}$ $10^{-5}$ 12.28 11.91 8.94 10.48 13.79 14.73 15.87 $10^{-4}$ 11.45 10.36 8.26 9.82 11.87 12.56 13.62 $10^{-3}$ 10.82 9.14 7.92 8.53 9.75 10.82 12.98 $5\times 10^{-3}$ 10.09 8.35 7.08 7.46 8.52 10.17 12.25 $10^{-2}$ 11.26 9.97 8.13 8.92 10.46 12.88 13.43 $5\times 10^{-2}$ 12.89 11.25 10.38 11.76 12.85 14,75 15.37 $10^{-1}$ 14.55 13.58 12.25 14.41 15.72 16.91 17.93
Appendix D Results of Comparison with Unstructured Pruning Methods
We compare ELDeR at a 25% pruning ratio with unstructured pruning methods employing 2:4 sparsity (where two out of every four consecutive parameters are set to zero) and present the experimental results in Table 8. For unstructured pruning methods, only calibration data is required. To ensure consistency, all methods utilize a total of 128 samples from WikiText-2 training set, randomly selected from WikiText-2. The models used in these experiments consists of the OPT and LLaMA model families.
The results are shown in Table 8. It can be seen that ELDeR with a pruning ratio of 25% significantly outperforms existing unstructured pruning methods with a sparsity of 2:4 on generative and zero-shot tasks, demonstrating the powerful performance of ELDeR. The reason ELDeR outperforms unstructured pruning in comparisons lies in iteratively learning layer weights and the regularization process before pruning. The regularization process transfers important information hidden in the parameter space which will be pruned to the remaining parts of the model, preventing irreversible information loss and performance degradation caused by direct pruning.
Table 8: Performance comparison with unstructured pruning methods. ‘PPL’ refers to the perplexity on WikiText-2. The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy. The best result is highlighted in bold, and the second-best in underlined. ELDeR-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while ELDeR-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
Model Method Pruning Ratio PPL ($\downarrow$) PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 Magnitude 2:4 269 56.42 41.79 34.12 31.38 16.19 35.98 Wanda 2:4 16.98 63.78 56.39 40.29 41.37 23.28 45.02 SparseGPT 2:4 13.96 66.52 54.71 46.48 45.72 24.89 47.66 ELDeR-$\ell_{2}$ 25% 13.18 70.54 60.27 46.35 51.59 27.36 51.22 ELDeR-$\ell_{1}$ 25% 13.12 70.65 60.13 46.24 51.76 27.55 51.27 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 Magnitude 2:4 978 61.98 47.26 42.19 40.53 20.18 42.43 Wanda 2:4 14.08 70.53 59.14 53.83 50.86 27.31 52.33 SparseGPT 2:4 12.25 72.18 59.74 56.26 52.34 29.11 53.93 ELDeR-$\ell_{2}$ 25% 11.26 73.15 62.26 60.58 55.93 30.37 56.46 ELDeR-$\ell_{1}$ 25% 11.49 72.89 63.09 60.23 56.21 30.16 56.52 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 Magnitude 2:4 1.21e4 62.37 50.28 47.57 42.26 20.67 44.63 Wanda 2:4 13.72 71.67 60.65 56.81 52.64 28.51 54.06 SparseGPT 2:4 12.89 73.65 61.08 58.47 55.36 30.58 55.83 ELDeR-$\ell_{2}$ 25% 10.45 74.15 64.47 68.82 61.55 35.23 60.84 ELDeR-$\ell_{1}$ 25% 10.32 74.58 64.37 68.45 61.29 34.87 60.71 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 Magnitude 2:4 56.75 63.21 59.75 55.41 58.27 26.32 52.59 Wanda 2:4 11.87 69.75 61.56 54.24 56.38 30.84 54.55 SparseGPT 2:4 9.62 71.45 63.82 57.34 59.81 33.75 57.23 ELDeR-$\ell_{2}$ 25% 7.08 72.48 67.52 60.45 62.69 39.73 60.57 ELDeR-$\ell_{1}$ 25% 7.17 72.08 67.93 60.42 62.38 39.89 60.54 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 Magnitude 2:4 8.75 69.38 61.56 49.43 61.82 30.65 54.57 Wanda 2:4 8.49 72.85 66.28 62.73 63.09 36.52 60.29 SparseGPT 2:4 7.15 74.72 67.68 64.23 65.19 38.58 62.08 ELDeR-$\ell_{2}$ 25% 5.82 74.56 69.34 64.79 71.25 45.63 65.11 ELDeR-$\ell_{1}$ 25% 5.89 75.06 69.18 64.96 71.43 45.26 65.18
Appendix E Performance of ELDeR under Different Pruning Ratios and Datasets
E.1 The Perplexity of ELDeR under Different Pruning Ratios and Datasets
To systematically evaluate the performance of ELDeR on zero-shot tasks across different pruning rates, we adopt the experimental setup outlined in Section 4.2, in which 128 samples are randomly sampled from the WikiText-2 training set to guide both the iterative learning of layer weights and the regularization process. Subsequently, we evaluate multiple large language models (LLMs) by measuring changes in perplexity across various generative task datasets, including WikiText-2, Alpaca, PTB, and C4, under pruning rates of 10%, 20%, 30%, 40%, 50%, and 60%. The detailed results, presented in Table 9, indicate that ELDeR exhibits greater robustness as model scale increases, suggesting that the proposed method effectively mitigates performance degradation in larger architectures. This highlights the scalability of ELDeR and its potential to maintain model efficiency under varying levels of sparsity.
Table 9: Perplexity comparison of ELDeR with different pruning ratios. We set the pruning ratios to 10%, 20%, 30%, 40%, 50%, and 60%, and test the perplexity of the OPT and LLaMA2 models on the generation task datasets Alpaca, WikiText-2, PTB, and C4. For ELDeR, we use the $\ell_{2}$-norm.
Model Pruning Ratio WikiText-2($\downarrow$) Alpaca($\downarrow$) PTB($\downarrow$) C4($\downarrow$) OPT-2.7B Dense 12.46 11.64 17.97 14.32 10% 12.78 11.89 18.65 15.02 20% 12.96 12.15 19.38 16.36 30% 15.52 13.47 24.28 20.75 40% 19.67 15.82 32.25 25.41 50% 25.62 21.89 47.67 33.73 60% 35.62 32.19 59.87 47.38 OPT-6.7B Dense 10.85 10.27 15.77 12.71 10% 11.03 10.95 16.78 13.46 20% 11.48 11.53 17.96 15.19 30% 12.87 12.75 20.86 17.52 40% 14.87 13.39 26.12 21.63 50% 21.43 16.51 35.36 28.74 60% 29.63 22.16 48.04 40.25 OPT-13B Dense 10.12 9.46 14.52 12.06 10% 10.28 9.89 15.12 12.73 20% 10.39 10.37 16.51 13.45 30% 11.38 11.12 19.42 16.14 40% 12.67 12.33 24.62 20.79 50% 15.38 14.76 30.88 27.65 60% 28.23 20.98 39.38 36.14 LLaMA2-7B Dense 5.47 5.25 7.92 7.26 10% 5.58 5.31 8.12 7.47 20% 6.13 5.87 8.78 7.92 30% 8.26 7.64 9.58 8.85 40% 10.28 9.79 12.87 11.42 50% 14.58 13.14 19.52 15.96 60% 25.18 21.46 30.14 28.32 LLaMA2-13B Dense 4.88 4.63 7.16 6.73 10% 4.99 4.91 7.48 7.93 20% 5.34 5.26 8.25 8.75 30% 6.87 6.35 8.97 9.68 40% 8.95 7.92 10.08 11.26 50% 11.23 9.73 12.63 14.73 60% 15.74 12.52 17.82 19.45
E.2 The Accuracy of ELDeR under Different Pruning Ratios on Zero-shot Tasks
To systematically evaluate the performance of ELDeR on zero-shot tasks across different pruning rates, we adopt the experimental setup outlined in Section 4.2, in which 128 samples are randomly sampled from the WikiText-2 training set to guide both the iterative learning of layer weights and the regularization process. We assess the accuracy of different model configurations at pruning rates of 10%, 20%, 30%, 40%, 50%, and 60% across a diverse set of benchmark datasets, including PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c. The results, summarized in Table 10, provide insights into the impact of sparsity on zero-shot generalization. Notably, the analysis reveals that ELDeR maintains competitive performance even at higher pruning rates, demonstrating its effectiveness in preserving reasoning and commonsense understanding across different tasks.
Table 10: Accuracy comparison of ELDeR with different pruning ratios. We set the pruning ratios to 10%, 20%, 30%, 40%, 50%, and 60%, and test the accuracy of the OPT and LLaMA2 models on the zero-shot task datasets PIQA, WinoGrande, HellaSwag, ARC-e and ARC-c. For ELDeR, we use the $\ell_{2}$-norm. ‘Avg_Acc’ represents the average accuracy.
Model Pruning Ratio PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) OPT-2.7B Dense 74.81 61.01 60.58 54.42 31.14 56.39 10% 72.35 60.75 51.53 52.73 29.15 53.30 20% 71.47 60.46 48.72 51.95 28.04 52.13 30% 66.73 58.53 44.36 50.02 26.58 49.24 40% 63.57 55.32 41.63 47.79 25.24 46.71 50% 58.48 52.29 40.52 46.26 21.32 43.77 60% 52.85 49.59 38.46 43.24 16.07 40.04 OPT-6.7B Dense 76.39 65.19 67.16 60.14 34.64 60.70 10% 75.42 63.38 65.16 57.23 32.89 58.82 20% 74.58 62.25 61.46 55.98 31.75 57.20 30% 71.88 60.56 59.15 53.62 28.54 54.75 40% 66.98 57.25 53.66 49.42 25.73 50.61 50% 62.78 54.56 45.63 46.38 21.75 46.22 60% 54.35 50.26 41.32 42.69 18.23 41.37 OPT-13B Dense 76.82 64.80 69.81 61.87 35.67 61.79 10% 75.49 64.67 69.26 61.73 35.54 61.34 20% 74.89 64.59 68.95 61.62 35.41 61.09 30% 71.46 62.67 66.53 59.39 33.61 58.73 40% 68.52 60.39 63.57 56.12 28.83 55.49 50% 63.35 57.62 57.43 51.65 24.09 50.83 60% 56.28 53.06 51.69 47.26 21.87 46.03 LLaMA2-7B Dense 79.11 69.06 75.99 74.58 46.25 69.00 10% 77.62 68.45 70.25 69.85 43.42 65.92 20% 75.38 67.96 65.26 65.83 41.57 63.20 30% 71.29 64.77 57.68 60.45 38.65 58.57 40% 67.61 60.38 54.12 57.53 35.45 55.02 50% 61.87 56.42 51.62 53.45 30.28 50.73 60% 55.39 51.54 48.73 47.42 27.08 46.03 LLaMA2-13B Dense 80.47 72.22 79.39 77.48 49.23 71.76 10% 78.45 71.87 74.32 76.26 48.39 69.86 20% 76.12 70.91 70.06 74.68 46.51 67.66 30% 72.54 68.15 63.51 70.88 43.21 63.66 40% 71.45 65.52 62.28 69.13 41.79 62.03 50% 67.36 58.52 60.65 63.45 37.88 57.57 60% 61.89 55.26 58.06 57.62 31.74 52.91
Appendix F Similarity Changes in Unregularized Layers
We plot the input-output similarity of the unregularized layers in LLaMA2-7B, as described in Section 4.9, in Figure 8. According to the results shown below, the similarity of all layers except for the first one is already high before regularization is applied. After applying regularization, the similarity in these unregularized layers decreases, indicating that the transformations undergone by the inputs in these layers become more substantial. This suggests that more information is being captured in these layers compared to before, implying that the regularization process may cause information to shift from the regularized layers to the unregularized ones.
Figure 8: A comparison between existing layer-wise structured pruning methods and ELDeR. Deeper blue layer represents greater performance impact, the taller cylinder represents larger data volume.
Appendix G Future Work and Limitations
ELDeR has primarily been evaluated on autoregressive language models and its applicability to other architectures or tasks remains unexplored. In future work, we plan to explore the application of ELDeR to other model architectures, such as convolutional neural networks (CNNs).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have provided the main claims of our research in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have provided the limitation and future work discussion in Section G.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Proposition 3.1 is theoretical. We provide its corresponding assumptions and proofs in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Please refer to Section 4.1 in the paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The dataset we used in this work is public available and we will release our reproducible code after the author notification.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Please refer to Section B in the paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: The reason error bars are not included is due to the high computational cost required to compute them.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided the experimental computing resources of our paper in Section 4.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have carefully reviewed the NeurIPS Code of Ethics and ensured full compliance, including preserving anonymity. Our study uses publicly available models and datasets, with proper citations to the original sources. No aspects of our work violate the ethical guidelines set by NeurIPS.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have provided the broader impacts of our work in Section 5 and Section G.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: There is no such risks in our work.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We ensure all papers and code used or relevant to our paper are properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We totally use public benchmarks.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We have described the usage of LLMs in our core method in Section 4.1.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
-ELDeR: Getting Efficient LLMs through Data-Driven Regularized Layer-wise Pruning
+Two-Stage Regularization-Based Structured Pruning for LLMs
 Abstract
-The deployment of Large language models (LLMs) in many fields is largely hindered by their high computational and memory costs.
-Recent studies suggest that LLMs exhibit sparsity, which can be used for pruning.
-Previous pruning methods typically follow a prune-then-finetune paradigm.
-Since the pruned parts still contain valuable information, statically removing them without updating the remaining parameters often results in irreversible performance degradation, requiring costly recovery fine-tuning (RFT) to maintain performance.
-To address this, we propose a novel paradigm: first apply regularization, then prune.
-Based on this paradigm, we propose ELDeR: Getting Efficient LLMs through Data-Driven Regularized Layer-wise Pruning.
-We multiply the output of each transformer layer by an initial weight, then we iteratively learn the weights of each transformer layer by using a small amount of data in a simple way.
-After that, we apply regularization to the difference between the output and input of the layers with smaller weights, forcing the information to be transferred to the remaining layers.
-Compared with direct pruning, ELDeR reduces the information loss caused by direct parameter removal, thus better preserving the model’s language modeling ability.
-Experimental results show that ELDeR achieves superior performance compared with powerful layer-wise structured pruning methods, while greatly reducing RFT computational costs.
-Since ELDeR is a layer-wise pruning method, its end-to-end acceleration effect is obvious, making it a promising technique for efficient LLMs.
+The deployment of large language models (LLMs) is largely hindered by their large number of parameters.
+Structural pruning has emerged as a promising solution.
+Prior structured pruning methods directly remove unimportant parameters based on certain metrics, which often causes knowledge loss and necessitates extensive retraining.
+To overcome this, we introduce a novel pruning method TRSP: Two-Stage Regularization-Based Structured Pruning for LLMs.
+Specifically, we multiply the output of each transformer layer by an initial learnable weight and iteratively learn these weights by adding their $\ell_{1}$-norm as a regularization term to the loss function, serving as the first-stage regularization.
+Subsequently, we apply additional regularization to the difference between the output and input of layers with smaller weights, encouraging the shift of knowledge to the preserved layers.
+This serves as the second-stage regularization.
+TRSP retains more knowledge and better preserves model performance than direct parameter elimination.
+Through extensive experimentation we show that TRSP outperforms strong layer-wise structured pruning methods without requiring retraining.
+As a layer-wise pruning method, it delivers notable end-to-end acceleration, making it a promising solution for efficient LLM deployment.
+Two-Stage Regularization-Based Structured Pruning for LLMs
+Mingkuan Feng1∗, Jinyang Wu1∗ Siyuan Liu2thanks: Equal Contribution., Shuai Zhang1, Ruihan Jin1, Feihu Che1, Pengpeng Shao1, Zhengqi Wen1, Jianhua Tao1 1Tsinghua University, 2Peking University
 1 Introduction
-Large language models (LLMs) have made remarkable progress in various tasks and fields, showcasing their strong capabilities [1, 2, 3, 4].
-As models grow larger, the increasing number of parameters results in substantial computational and memory demands, posing major challenges to the practical deployment of LLMs.
-There is an urgent need for techniques that can compress model size without compromising performance.
-The main methods for reducing model size include pruning [5, 6, 7], quantization [8, 9], distillation [10], and low-rank decomposition [11].
-This study primarily concentrates on pruning, an effective and widely adaptable technique that is orthogonal to other model compression methods.
-Pruning methods are typically divided into two main types: unstructured pruning [12, 13] and structured pruning [14, 15].
-In contrast to unstructured pruning, structured pruning enables recovery fine-tuning (RFT) for specific downstream tasks without the need for specialized hardware.
-It can also achieve significant end-to-end acceleration without the need for specially designed hardware [16].
-Despite advancements, existing structured pruning methods have some limitations.
-They all follow the paradigm of first prune channels or layers based on a designed metric, and then performing RFT.
-However, they overlook that important information may also exist in the pruned parts [17, 18, 19], and directly removing them without updating the remaining parameters often results an irreversible decline in model performance.
-Therefore, obtaining satisfactory performance with pruned models typically requires a large amount of data for RFT [14], increasing computational overhead.
-Moreover, high sparsity can cause the model’s performance to collapse, limiting the acceleration effect.
-To address these limitations, we present a new pruning paradigm that first apply regularization and followed by pruning.
-The regularization process dynamically transfers crucial information from the layers to be pruned to the remaining layers, thus minimizing the information loss resulting from direct parameter removal.
-This paradigm can maintain good model performance without RFT.
-Based on this paradigm, we propose ELDeR, a simple and effective data-driven regularized layer-wise pruning method.
-The comparison of the ELDeR with existing layer-wise pruning approaches is shown in Figure 1.
-Specifically, the process of ELDeR is divided into the following steps: (1) select a small amount of data for learning layer weights and regularization, (2) iteratively learn the layer weights, (3) apply regularization to the difference between the outputs and inputs of the layers to be pruned, (4) prune.
-First, we randomly choose a small subset of data from commonly used benchmark datasets.
-Since the selected data is small, the computational costs in learning the layer weights and regularization process can be greatly minimized.
-Second, we iteratively learn the weights of each layer by adding their $\ell_{1}$-norm in the loss function.
-Third, we apply regularization ($\ell_{1}$-norm or $\ell_{2}$-norm) to the difference between the outputs and inputs of the layers with smaller weights, forcing important information to be transferred to the remaining layers which significantly reduces the performance decline caused by parameter removal.
-Thus the model can maintain good performance even at high pruning ratios, thereby providing significant end-to-end acceleration.
+Large language models (LLMs) have made remarkable progress in natural language processing Yang et al.
+(2024a); Wu et al.
+(2024); Guo et al.
+(2025).
+However, their large scale makes real-world deployment challenging.
+There is an urgent need for techniques that can enhance the compactness and computational efficiency of LLMs while preserving their language modeling capabilities.
+Structured pruning is a method used to simplify neural networks by removing unnecessary or redundant parameters Xia et al.
+(2024); An et al.
+(2024); Feng et al.
+(2025).
+Structured pruning is categorized into channel-wise pruning Ma et al.
+(2023); Ashkboos et al.
+(2024) and layer-wise pruning Song et al.
+(2024).
+Channel-wise pruning operates at the row or column level of parameter matrices.
+Layer-wise pruning operates at the level of entire transformer layers thereby offering a more simple approach compared to channel-wise pruning Chen et al.
+(2024); Men et al.
+(2024).
+However, existing layer-wise pruning methods have a certain limitation.
+They consistently first compute the importance of each transformer layer using a designed criteria, prune unimportant layers, and then fine-tune the pruned model to compensate for performance degradation caused by pruning.
+However, even unimportant layers can hold valuable knowledge Dettmers et al.
+(2022); Yin et al.
+(2024); An et al.
+(2025).
+This sequential process of selecting and then directly pruning layers does not handle the important knowledge contained in the layers that are to be pruned, resulting in its direct loss.
+The performance drop requires substantial retraining for recovery, leading to considerable computational overhead Ma et al.
+(2023).
+To address this, we present TRSP that first apply two-stage regularization and then prune.
+The first regularization process iteratively learns layer weights.
+The second regularization process dynamically transfers valuable knowledge from the layers to be pruned to the remaining layers in advance, greatly reducing the knowledge loss caused by pruning.
+The comparison of TRSP with existing layer-wise pruning approaches is shown in Figure 1.
+First, we sample a small portion of data from standard benchmark datasets randomly.
+Given the limited scale of the selected data, the computational overhead incurred during the two-stage regularization is significantly reduced.
+Second, we multiply the output of each transformer layer by an initial learnable weight and iteratively learn these weights by incorporating their $\ell_{1}$-norm as a regularization term in the loss function.
+This serves as the first stage regularization.
+Third, we apply regularization ($\ell_{1}$-norm or $\ell_{2}$-norm) to the difference between the outputs and inputs of the layers with smaller weights, forcing important knowledge to be transferred to the remaining layers which significantly reduces the performance decline caused by parameter removal.
+Thus the model can maintain good language modeling capability.
+This serves as the second stage regularization.
 Finally, we prune the layers with smaller weights.
-Extensive experiments demonstrate that ELDeR substantially outperforms current layer-wise pruning methods in terms of perplexity and accuracy across different pruning ratios, while also significantly improving end-to-end acceleration.
-The main contributions of ELDeR are summarized as follows:
-  • Propose A New Paradigm: ELDeR reduces information loss caused by direct parameter removal by sequentially learning layer weights, applying regularization ($\ell_{1}$-norm or $\ell_{2}$-norm), and performing pruning.
-This approach helps preserve model performance without requiring for RFT.
-  • High Performance: ELDeR outperforms previous layer-wise structured pruning methods in both generation and zero-shot tasks.
-The pruned model demonstrates a considerable acceleration compared to the dense model.
-  • Low Overhead: By using only a small amount of data for learning layer weights and regularization, ELDeR significantly reduces the overhead of RFT.
-Figure 1: A comparison between existing layer-wise structured pruning methods and ELDeR. Deeper blue layer represents greater performance impact, the taller cylinder represents larger data volume.
+Comprehensive experiments demonstrate that TRSP substantially outperforms strong layer-wise pruning methods in generation tasks and zero-shot tasks across different pruning ratios, while also significantly improving end-to-end acceleration.
+The main contributions of TRSP are summarized as follows:
+  • Retention of Knowledge: TRSP reduces knowledge loss by progressively applying two-stage regularization and performing pruning.
+This approach helps preserve model performance without requiring for retraining.
+  • Effectiveness: TRSP outperforms strong layer-wise pruning methods in generation and zero-shot tasks. The pruned model demonstrates a considerable acceleration.
+  • Minimal Cost: The data required for two-stage regularization is minimal and TRSP is retraining free after pruning.
+Figure 1: A comparison between existing layer-wise structured pruning methods and TRSP . Deeper blue layer represents greater performance impact, the taller cylinder represents larger data volume.
 2 Related Works
-2.1 Pruning Methods
-The process of removing unnecessary weights, known as pruning, has proven to be an effective method for reducing the complexity of deep neural networks over the past few decades [20, 21, 22, 23].
-Pruning methods can generally be divided into two categories: unstructured pruning [24, 25, 26] and structured pruning [5, 27, 28].
-Unstructured pruning methods eliminate individual weights according to their significance.
-For instance, Magnitude [22, 29] assumes larger weights are more important, Wanda [13] uses the product of the layer weights and activations as the criterion, and SparseGPT [12] uses the weight squared divided by the inverse Hessian matrix.
+2.1 Model Pruning
+Model Pruning aims to improve model efficiency by sparsification or parameter removal LeCun et al.
+(1989); Hassibi et al.
+(1993); Han et al.
+(2015); Liu et al.
+(2017).
+Several studies employ unstructured Kurtic et al.
+(2022); Zhang et al.
+(2024); Xu et al.
+(2024) and structured pruning Xia et al.
+(2024); Yang et al.
+(2024b); Gao et al.
+(2024b).
+Unstructured pruning zeros individual neurons according to their importance such as SparseGPT Frantar and Alistarh (2023), SpQR Dettmers et al.
+(2024), Pruner-Zero Dong et al.
+(2024), and Wanda Sun et al.
+(2024) .
 Their main advantage is flexibility.
-However, their drawbacks include the need for specialized hardware to accelerate [30], the requirement for extreme sparsity to achieve considerable acceleration [31], and are not suitable for RFT on downstream tasks.
-Structured pruning methods, including channel-wise pruning [15] and layer-wise pruning [32].
-Channel-wise pruning methods typically create a metric to assess the importance of channels in the parameter matrixs of LLMs, and then remove the less important ones.
+However, they need dedicated hardware to accelerate Xia et al.
+(2023), and are not able to retrain on downstream tasks.
+Structured pruning methods can be categorized by granularity into channel-wise pruning Ashkboos et al.
+(2024) and layer-wise pruning Kim et al.
+(2024a).
+Channel-wise pruning methods create a metric to assess the significance of channels in the parameter matrices of LLMs, and then remove the less significant ones.
 Layer-wise pruning methods treat entire layers as the basic units for pruning.
-For instance, SLEB [16], iteratively remove entire transformer layers by evaluating their impact on the model’s final loss, ShortGPT [33] thinks high similarity between layers means redundancy, LaCo [27] uses layer collapse to prune, Shortened LLaMA [34] prune layers in one-shot based on their importance.
-However, these methods have a major limitation: even less important channels or layers may still contain valuable information, and directly pruning them often results in significant performance degradation.
-To overcome this, we explore the possibility of "shifting" important information from the pruned parts to the remaining parts, which could greatly maintain model performance.
+For instance, SLEB Song et al.
+(2024) iteratively remove entire transformer layers by evaluating their impact on the model’s final loss, ShortGPT Men et al.
+(2024) thinks high similarity between layers means redundancy, LaCo Yang et al.
+(2024b) uses layer collapse to prune, Shortened LLaMA Kim et al.
+(2024b) prune layers in one-shot based on their importance.
+In this paper, we focus on layer-wise pruning.
+Prior layer-wise pruning approaches suffer from the drawback that the less important layers may still carry critical knowledge, and pruning them often causes knowledge loss.
+Discovering a way to reshape knowledge distribution prior to pruning could potentially alleviate knowledge loss.
 2.2 Regularization
-Regularization is commonly used in machine learning [35, 36, 37], including preventing model overfitting [38] and feature selection [39].
-The $\ell_{1}$-norm typically promotes sparsity by driving certain coefficients to zero, while the $\ell_{2}$-norm encourages smoother solutions [40].
-Both can significantly affect the distribution pattern of the data [22, 41].
-Inspired by this, we propose a new pruning paradigm, in which the regularization process could transfer important information from the pruned parameter space to the remaining parts of the model, thus improving performance.
+In machine learning, regularization plays a vital role in controlling overfitting Santos and Papa (2022) and identifying informative features Tibshirani (1996), and has been extensively studied Hoerl and Kennard (1970); Poggio et al.
+(1987); Balestriero et al.
+(2022).
+The $\ell_{1}$-norm tends to enforce compact representations by eliminating certain parameters, whereas the $\ell_{2}$-norm favors stability and continuity Boyd and Vandenberghe (2004).
+Both can both alter the underlying structure and representation of the data Han et al.
+(2015); Tao et al.
+(2023).
+Inspired by this insight, regularization can be leveraged to migrate critical knowledge from pruned layers to preserved layers, thereby enhancing model performance.
 3 Methodology
-In this section, we provide a detailed introduction to ELDeR. The overall ELDeR algorithm is summarized in Algorithm 1. ELDeR mainly consists of four steps:
-  • Select Data: Select a small amount of data to learn the layer weights and apply regularization.
-  • Learn Layer Weights: Iteratively learn the weights of each layer by adding their $\ell_{1}$-norm to the loss function.
-  • Regularization: Apply regularization to the difference between the output and input of layers with smaller weights, facilitating information transfer.
-  • Pruning: Removes the layers that were regularized in the previous step.
-3.1 Select Data
-The data is selected from widely used datasets, including Alpaca [42], WikiText-2 [43], PTB [44], and C4 [45].
-For instance, approximately 128 samples are randomly selected from the WikiText-2 training dataset for learning layer weights and applying regularization.
-3.2 Learn Layer Weights
-For clarity, we first define the relevant notations.
-$\mathbf{W}$ represents the initial model.
-Let $p$ be the pruning ratio, indicating that $p\%$ of model parameters will be pruned.
-The number of layers in the initial model is denoted by $l$.
-The model hidden size is $d$.
-Let $\mathbf{X}\in\mathbb{R}^{b\times n\times d}$ represents the input data embedding, where $b$ is the batch size and $n$ is the number of tokens.
-The input to the $i$th layer is denoted as $\mathbf{X_{\text{in}}^{i}}\in\mathbb{R}^{b\times n\times d}$, and the output from the $i$th layer is denoted as $\mathbf{X_{\text{out}}^{i}}\in\mathbb{R}^{b\times n\times d}$.
-A learnable weight $S[i]$ is assigned to the $i$th transformer layer, and the set of all layer weights is denoted as $S$.
-Algorithm 1 ELDeR algorithm.
-1:Input: selected data: $\mathbf{X}$, number of layers in LLM: $l$, initial model $\mathbf{W}$, number of layers to prune: $n$, set of layer weights: $S$, set of layers to be pruned: $P$, norm type: flag.
+The complete TRSP procedure is outlined in Algorithm 1 and Figure 1.
+The TRSP framework involves four key stages: (1) Prepare data: Select a small amount of data for the following two-stage regularization.
+(2) Learn layer weights: Iteratively learn the weights of each layer by incorporating their $\ell_{1}$-norm as a regularization term in the loss function.
+(3) The second stage regularization: Apply regularization to the difference between the output and input of layers with smaller weights, facilitating knowledge transfer.
+(4) Pruning: Removes the layers that were regularized in the previous step.
+Algorithm 1 TRSP algorithm.
+1:Input: selected data: $\mathbf{X}$, number of layers: $l$, initial model $\mathbf{W}$, number of layers to prune: $n$, layer weights: $S$, set of pruned layers: $P$, norm type: flag.
 2:Initialize each layer weight in $S$ to 1.
@@
 20:$\mathbf{W}\leftarrow Prune(\mathbf{W},P)$
+3.1 Prepare Data
+The data is drawn from standard benchmarks, including Alpaca Taori et al.
+(2023), WikiText-2 Merity et al.
+(2016), PTB Marcus et al.
+(1993), and C4 Raffel et al.
+(2020).
+For example, 128 instances are randomly drawn from the WikiText-2 training set to for layer weight learning via regularization and the second stage regularization.
+3.2 Learn Layer Weights
+This is the first stage regularization.
+To better understand our paper, we first define some notations.
+$\mathbf{W}$ represents the initial model.
+Let $p$ be the pruning ratio, indicating that $p\%$ of model layers will be pruned.
+The number of layers in the initial model is $l$.
+The model hidden size is $d$.
+Let $\mathbf{X}\in\mathbb{R}^{b\times n\times d}$ represents the data embedding, where $b$ is the batch size and $n$ is the number of tokens.
+The input to the $i$th layer is denoted as $\mathbf{X_{\text{in}}^{i}}\in\mathbb{R}^{b\times n\times d}$, and the output from the $i$th layer is denoted as $\mathbf{X_{\text{out}}^{i}}\in\mathbb{R}^{b\times n\times d}$.
+A learnable weight $S[i]$ is assigned to the $i$th transformer layer, and the set of all layer weights is denoted as $S$.
 According to Algorithm 1, we initialize the weight of each layer to 1.
-As shown in Figure 2, the output of each layer is scaled by its associated weight before being passed as input to the subsequent layer.
+As shown in Figure 2, the output of each layer is scaled by its associated weight before being passed as input to the next layer.
 To learn the weight of each layer, we employ the input data embedding $\mathbf{X}$.
 When pruning $n$ layers, $\mathbf{X}$ is repeatedly used as input in each iteration.
-The objective function defined in Equation 1 comprises two parts: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$, and the sum of the $\ell_{1}$-norm of all layer weights, $\lambda_{1}$ balances the two components, $l_{1}$ is the number of layers in the current model that are not masked.
+The objective function in Equation 1 comprises two parts: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$, and the sum of the $\ell_{1}$-norm of all layer weights, $\lambda_{1}$ balances the two components, $l_{1}$ is the number of layers in the current model that are not masked.
 $\begin{split}\mathcal{L}_{\text{learn}}=\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{1}\sum_{i=0}^{l_{1}-1}\|S[i]\|_{1}\end{split}$ (1)
@@
 However, as shown in Section 4.8, this often results in the removal of consecutive transformer layers, which leads to a substantial degradation in model performance.
-We formalize the minimization of the objective function in Equation 1 as an optimization problem.
+The process of minimizing the function in Equation 1 is treated as an optimization task.
 Since $\ell_{1}$-norm is not differentiable, backpropagation (BP) can’t be used directly, we need to transform the problem using Proposition 3.1.
-Proposition 3.1.
-(If the loss function includes the $\ell_{1}$-norm, it can also be solved using BP.
+Proposition 3.1
+(If the objective function contains an $\ell_{1}$ regularization term, it can still be optimized using BP.
 Proof in Appendix A).
 The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
-$\min\quad||x||_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (2)
+$\min\quad||x||_{1}\hskip 5.0pt\iff\hskip 5.0pt\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (2)
 s.t. $\quad-y\leq x\leq y,$
 $\quad y\geq 0.$
-The optimization problem in Equation 1 can be equivalently transformed into a constrained formulation in Equation 3.
-After transformation, the objective function is differentiable, and can be solved by BP.
+The objective function in Equation 1 can be reformulated as an equivalent constrained problem in Equation 3. After transformation, the objective function is differentiable, and can be solved by BP.
 $\min_{W,y}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{1}\mathbf{1}^{T}y$ (3)
@@
 $\quad y\geq 0.$
-3.3 Regularization
-After completing learning the layer weighs, we further lighten the impact of the pruned layers on the
+3.3 The Second Stage Regularization
 Figure 2: Details of layer weight learning and regularization. $P$ is the set of prunable layers, identified iteratively using the method in Section 3.2.
-final model output by applying regularization to them.
-The approach is simple and straightforward: we use the input data embedding $\mathbf{X}$ as input and apply a one-shot regularization on the difference between the output and input of each layer in the pruning set $P$.
+After learning the layer weighs, we further lighten the impact of the pruned layers on the final model output by applying regularization to them.
+The approach is simple and straightforward: we use the data embedding $\mathbf{X}$ as input and apply a one-shot regularization on the difference between the output and input of each layer in the pruning set $P$.
 For example, in Figure 2, the $i$th layer in $P$ and should be regularized, we just add the $\ell_{1}$-norm or $\ell_{2}$-norm of $\mathbf{X_{out}^{i}-X_{in}^{i}}$ to the loss function.
-This encourages the information to be redistributed from the pruned layers to the remaining ones, thereby reducing the amount of information retained in the pruned layers and significantly minimizing the performance degradation caused by subsequent pruning.
-The objective function defined in Equation 4 comprises two components: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$ and the regularization loss, $\lambda_{2}$ balances the two components.
+This encourages the knowledge to be redistributed from the pruned layers to the remaining ones, thereby reducing the amount of knowledge retained in the pruned layers and significantly minimizing the performance degradation caused by subsequent pruning.
+The objective function in Equation 4 comprises two parts: the language modeling loss $\mathcal{L}(\mathbf{W},\mathbf{X})$ and the regularization loss, $\lambda_{2}$ balances the two components.
 When the $\ell_{1}$-norm is used, the equivalent constrained optimization problem transformed using Proposition 3.1 is shown in Equation 5.
@@
 $\begin{split}\mathcal{L}_{\text{sum}}=\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{2}\sum_{i=0}^{|P|-1}\|\mathbf{X_{out}^{i}-X_{in}^{i}}\|\end{split}$ (4)
-$\min_{W,Y_{i}}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{2}\sum_{i=0}^{|P|-1}\mathbf{1}^{T}\mathbf{Y_{i}}\mathbf{1}^{T}$ (5)
+$\min_{W,Y_{i}}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda_{2}\sum_{i=0}^{|P|-1}\mathbf{1}^{T}\mathbf{Y_{i}}\mathbf{1}$ (5)
 s.t. $\quad-\mathbf{Y_{i}}\leq\mathbf{X_{\text{out}}^{i}-X_{\text{in}}^{i}}\leq\mathbf{Y_{i}},$
@@
 3.4 Pruning
-After applying the regularization, we directly remove the transformer layers in the set $P$, which is a straightforward process.
+After applying the regularization, we directly remove the transformer layers in the set $P$.
 4 Experiments
-This section introduces experimental setup (4.1) and analyzes the effectiveness of ELDeR from the following aspects: performance comparison (4.2), acceleration 4.3, robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (4.7), choice of learning layer weights (4.8), impact of regularization (4.9).
+This section introduces experimental setup (4.1) and analyzes the effectiveness of TRSP from the following aspects: performance comparison (4.2), acceleration 4.3, robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (3), choice of learning layer weights (4.8), impact of regularization (4.9).
 4.1 Experimental Setup
-Implementation: All methods are implemented in PyTorch [46], using the Hugging Face Transformers library [47].
-All experiments are conducted on 80GB NVIDIA A100 GPUs.
-For fairness, we use llm-eval-harness [48] to evaluate the pruned models.
+Datasets: We evaluated on generation and zero-shot tasks.
+For generation task, following prior workAshkboos et al.
+(2024), we evaluate the model’s perplexity on WikiText-2 Merity et al.
+(2016) test set.
+For zero-shot task, we evaluate on PIQA Bisk et al.
+(2020), WinoGrande Sakaguchi et al.
+(2021), HellaSwag Zellers et al.
+(2019), ARC-e and ARC-c Clark et al.
+(2018).
+In Section 4.5 we randomly selected data from Alpaca Taori et al.
+(2023), WikiText-2 Merity et al.
+(2016), PTB Marcus et al.
+(1993), and C4 Raffel et al.
+(2020).
+Implementation: All methods are developed in PyTorch Paszke et al.
+(2019), leveraging the Hugging Face Transformers library Wolf (2019).
+Experimental evaluations are carried out on 80GB NVIDIA A100 GPUs.
+Pruned models are evaluated with llm-eval-harness Gao et al.
+(2024a).
 More details are in Appendix B.
-Datasets: We evaluated on generation and zero-shot tasks.
-For generation task, following prior work[15], we evaluate the model’s perplexity on WikiText-2 [43] test set.
-For zero-shot task, we evaluate on PIQA [49], WinoGrande [50], HellaSwag [51], ARC-e and ARC-c [52].
-In Section 4.5 we use a randomly selected subset of data from Alpaca [42], WikiText-2 [43], PTB [44], and C4 [45].
-Models: The models pruned using ELDeR include the LLaMA models (LLaMA2-7B, LLaMA2-13B, LLaMA3-8B) [1, 53], OPT models (OPT-2.7B, OPT-13B) [54], and Phi-2 [55].
-Baselines: We evaluate ELDeR against competitive layer-wise structured pruning methods: SLEB [16], ShortGPT [33], LaCo [27], Shortened LLaMA (PPL version) [34].
-We also compared with powerful unstructured pruning methods like Magnitude [22], Wanda [13], and SparseGPT [12] in Appendix D.
-Evaluation Metrics: The performance on generation task is measured by perplexity, a well-established and robust metric [56], while zero-shot tasks performance is evaluated using accuracy [57].
-The acceleration effects are represented by throughput and latency [16].
+Evaluation Metrics: The ability on generation task is evaluated by perplexity, a well-established and robust metric Yao et al.
+(2022).
+Zero-shot tasks ability is evaluated using accuracy Dong et al.
+(2024).
+Acceleration is measured by throughput and latency Song et al.
+(2024).
+Models: The models include the Phi-2 Javaheripi et al.
+(2023), OPT models (OPT-2.7B, OPT-13B) Zhang et al.
+(2022), and LLaMA models (LLaMA2-7B, LLaMA2-13B, LLaMA3-8B) Touvron et al.
+(2023); Grattafiori et al.
+(2024).
+Baselines: We compare TRSP with strong layer-wise structured pruning methods: SLEB Song et al.
+(2024), ShortGPT Men et al.
+(2024), LaCo Yang et al.
+(2024b), Shortened LLaMA (PPL version) Kim et al.
+(2024b).
 4.2 Performance Comparison
-For fairness, we randomly selected 128 samples from the WikiText-2 training dataset with a sequence length of 2048 as the data for ELDeR to get layer weights and perform regularization, as well as the calibration data for the baselines.
-Following [15], we selected another 1,000 samples from WikiText-2 training dataset to perform RFT leveraging LoRA [58] on the baselines after pruning.
+To ensure fairness, 128 sequences of length 2048 were randomly drawn from the WikiText-2 training set for TRSP’s two-stage regularization and the calibration process of the baselines.
+Following Ashkboos et al.
+(2024), we selected another 1,000 samples from WikiText-2 training dataset to retrain leveraging LoRA Hu et al.
+(2022) on the baselines after pruning.
 The aforementioned 128 samples and the 1,000 samples have no overlap.
-Because ELDeR is finetuning-free, we did not perform RFT on ELDeR.
-The pruning ratio was 25% which means 25% of the transformer layers in a model will be pruned.
-Table 1: Performance comparison of different pruning methods.
-‘PPL’ refers to perplexity on WikiText-2.
-The accuracy is reported on five zero-shot benchmarks.
-The best result is in bold, and the second-best is underlined.
-ELDeR-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, ELDeR-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
+Because TRSP is retraining-free, we did not retrain TRSP.
+The pruning ratio was 25% which means 25% of the transformer layers in a model will be removed.
+Table 1: Performance comparison of TRSP and baselines.
 ‘PR’ is the pruning ratio.
-Model Method PR PPL ($\downarrow$) PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 SLEB 25% 7.65 68.85 63.63 50.96 49.38 30.79 52.72 ShortGPT 25% 7.15 69.67 65.19 51.26 52.46 33.89 54.49 LaCo 25% 7.38 68.53 63.76 50.39 51.28 33.45 53.48 Shortened LLaMA 25% 7.74 66.88 62.19 51.45 51.47 32.66 52.93 ELDeR-$\ell_{2}$ 25% 6.53 71.35 67.62 55.84 52.26 35.75 56.56 ELDeR-$\ell_{1}$ 25% 6.58 71.42 67.13 56.05 52.18 35.79 56.51 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 SLEB 25% 15.71 65.19 56.26 44.54 46.28 25.14 47.48 ShortGPT 25% 14.96 67.37 57.65 46.82 49.43 26.54 49.56 LaCo 25% 15.38 66.54 59.45 43.68 48.74 24.96 48.67 Shortened LLaMA 25% 15.89 63.14 57.36 43.57 47.62 25.88 47.51 ELDeR-$\ell_{2}$ 25% 13.18 70.54 60.27 46.35 51.59 27.36 51.22 ELDeR-$\ell_{1}$ 25% 13.12 70.65 60.13 46.24 51.76 27.55 51.27 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 SLEB 25% 9.63 65.22 63.38 55.51 56.39 33.46 54.79 ShortGPT 25% 8.89 66.75 66.26 57.14 58.93 36.42 57.10 LaCo 25% 9.14 69.45 65.31 52.67 55.73 34.89 55.61 Shortened LLaMA 25% 9.47 65.58 64.72 58.36 54.19 32.96 55.16 ELDeR-$\ell_{2}$ 25% 7.08 72.48 67.52 60.45 62.69 39.73 60.57 ELDeR-$\ell_{1}$ 25% 7.17 72.08 67.93 60.42 62.38 39.89 60.54 LLaMA3-8B Dense 0% 5.76 85.56 77.94 79.27 78.84 56.49 75.62 SLEB 25% 10.38 72.74 64.12 67.74 65.84 45.16 63.12 ShortGPT 25% 9.26 75.38 69.25 70.12 68.54 47.57 66.17 LaCo 25% 10.14 74.69 67.52 66.36 69.43 46.31 64.86 Shortened LLaMA 25% 9.84 73.62 68.73 68.55 66.16 43.39 64.09 ELDeR-$\ell_{2}$ 25% 7.84 77.25 71.63 72.26 71.49 49.58 68.44 ELDeR-$\ell_{1}$ 25% 7.68 77.36 71.33 72.82 70.75 49.66 68.38 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 SLEB 25% 11.96 72.83 64.06 63.32 59.98 34.65 58.97 ShortGPT 25% 11.38 73.59 64.52 65.68 60.41 34.99 59.84 LaCo 25% 11.79 73.96 63.24 62.17 61.46 33.28 58.82 Shortened LLaMA 25% 11.62 71.26 63.57 66.29 58.47 33.83 58.68 ELDeR-$\ell_{2}$ 25% 10.45 74.15 64.47 68.82 61.55 35.23 60.84 ELDeR-$\ell_{1}$ 25% 10.32 74.58 64.37 68.45 61.29 34.87 60.71 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 SLEB 25% 7.08 68.31 66.86 57.12 62.19 38.45 58.59 ShortGPT 25% 6.79 73.26 68.37 62.25 67.54 43.38 62.96 LaCo 25% 7.14 72.35 65.78 59.43 65.36 42.73 61.13 Shortened LLaMA 25% 6.92 69.45 66.38 59.86 65.25 39.12 60.01 ELDeR-$\ell_{2}$ 25% 5.82 74.56 69.34 64.79 71.25 45.63 65.11 ELDeR-$\ell_{1}$ 25% 5.89 75.06 69.18 64.96 71.43 45.26 65.18
-As shown in Table 1, ELDeR achieves the lowest perplexity and the highest average accuracy across all models compared to the baseline methods, demonstrating its superior performance on both generation and zero-shot tasks.
+‘PPL’ is the perplexity on WikiText-2.
+The accuracy is evaluated on five zero-shot benchmarks.
+TRSP -$\ell_{2}$ means using the $\ell_{2}$-norm in the second stage regularization, TRSP -$\ell_{1}$ is using the $\ell_{1}$-norm.
+The best result is in bold, the second-best is underlined.
+Model Method PR PPL ($\downarrow$) PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 SLEB 25% 7.65 68.85 63.63 50.96 49.38 30.79 52.72 ShortGPT 25% 7.15 69.67 65.19 51.26 52.46 33.89 54.49 LaCo 25% 7.38 68.53 63.76 50.39 51.28 33.45 53.48 Shortened LLaMA 25% 7.74 66.88 62.19 51.45 51.47 32.66 52.93 TRSP -$\ell_{2}$ 25% 6.53 71.35 67.62 55.84 52.26 35.75 56.56 TRSP -$\ell_{1}$ 25% 6.58 71.42 67.13 56.05 52.18 35.79 56.51 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 SLEB 25% 15.71 65.19 56.26 44.54 46.28 25.14 47.48 ShortGPT 25% 14.96 67.37 57.65 46.82 49.43 26.54 49.56 LaCo 25% 15.38 66.54 59.45 43.68 48.74 24.96 48.67 Shortened LLaMA 25% 15.89 63.14 57.36 43.57 47.62 25.88 47.51 TRSP -$\ell_{2}$ 25% 13.18 70.54 60.27 46.35 51.59 27.36 51.22 TRSP -$\ell_{1}$ 25% 13.12 70.65 60.13 46.24 51.76 27.55 51.27 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 SLEB 25% 9.63 65.22 63.38 55.51 56.39 33.46 54.79 ShortGPT 25% 8.89 66.75 66.26 57.14 58.93 36.42 57.10 LaCo 25% 9.14 69.45 65.31 52.67 55.73 34.89 55.61 Shortened LLaMA 25% 9.47 65.58 64.72 58.36 54.19 32.96 55.16 TRSP -$\ell_{2}$ 25% 7.08 72.48 67.52 60.45 62.69 39.73 60.57 TRSP -$\ell_{1}$ 25% 7.17 72.08 67.93 60.42 62.38 39.89 60.54 LLaMA3-8B Dense 0% 5.76 85.56 77.94 79.27 78.84 56.49 75.62 SLEB 25% 10.38 72.74 64.12 67.74 65.84 45.16 63.12 ShortGPT 25% 9.26 75.38 69.25 70.12 68.54 47.57 66.17 LaCo 25% 10.14 74.69 67.52 66.36 69.43 46.31 64.86 Shortened LLaMA 25% 9.84 73.62 68.73 68.55 66.16 43.39 64.09 TRSP -$\ell_{2}$ 25% 7.84 77.25 71.63 72.26 71.49 49.58 68.44 TRSP -$\ell_{1}$ 25% 7.68 77.36 71.33 72.82 70.75 49.66 68.38 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 SLEB 25% 11.96 72.83 64.06 63.32 59.98 34.65 58.97 ShortGPT 25% 11.38 73.59 64.52 65.68 60.41 34.99 59.84 LaCo 25% 11.79 73.96 63.24 62.17 61.46 33.28 58.82 Shortened LLaMA 25% 11.62 71.26 63.57 66.29 58.47 33.83 58.68 TRSP -$\ell_{2}$ 25% 10.45 74.15 64.47 68.82 61.55 35.23 60.84 TRSP -$\ell_{1}$ 25% 10.32 74.58 64.37 68.45 61.29 34.87 60.71 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 SLEB 25% 7.08 68.31 66.86 57.12 62.19 38.45 58.59 ShortGPT 25% 6.79 73.26 68.37 62.25 67.54 43.38 62.96 LaCo 25% 7.14 72.35 65.78 59.43 65.36 42.73 61.13 Shortened LLaMA 25% 6.92 69.45 66.38 59.86 65.25 39.12 60.01 TRSP -$\ell_{2}$ 25% 5.82 74.56 69.34 64.79 71.25 45.63 65.11 TRSP -$\ell_{1}$ 25% 5.89 75.06 69.18 64.96 71.43 45.26 65.18
+As shown in Table 1, TRSP achieves the lowest perplexity and the highest average accuracy across all models, demonstrating its superior performance on both generation and zero-shot tasks.
+Notably, on OPT-13B, TRSP only drop 1% in average accuracy compared to the dense model.
 On LLaMA2-7B, its perplexity is 20% lower than the second-best method, ShortGPT.
-Notably, on OPT-13B, compared to the dense model, ELDeR only drop 1% in average accuracy.
-This further demonstrates the effectiveness of ELDeR.
-The performance difference between $\ell_{2}$-norm and $\ell_{1}$-norm is tiny, indicating that ELDeR is insensitive to the choice of regularization norm.
-In the following sections, we refer to ELDeR as the version that uses the $\ell_{2}$-norm.
+This further demonstrates the effectiveness of TRSP.
+The minimal performance difference between $\ell_{2}$-norm and $\ell_{1}$-norm suggests that TRSP is not sensitive to the choice of regularization norm.
+In the following sections, we refer to using the $\ell_{2}$-norm in the stage two regularization.
 4.3 Acceleration
-Language processing in LLMs consists of two main stages: prompt processing, which is compute-bound, and token generation, which is memory-bound.
-We analyze the speedup achieved in each stage separately.
-Table 2 shows the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated on a single 80GB NVIDIA A100 GPU.
-Following the methodology of previous work [16], the token generation test involves generating 128-token sentences with a batch size of 64, while prompt processing latency is measured using an input sequence of 2048 tokens.
-Table 2: Throughput (tokens/s) and latency (ms) on OPT-13B and LLaMA2-13B. ‘PPL’ refers to the perplexity on Wikitext2. ‘TI’ is Throughput Increase.
-Model Method Pruning ratio PPL($\downarrow$) Avg_Acc(%) Tokens/s($\uparrow$) TI($\uparrow$) Latency($\downarrow$) Speedup($\uparrow$) OPT-13B Dense 0% 10.12 61.79 1029 1.00× 386.5 1.00× ELDeR 25% 10.45 60.84 1348 1.31× 286.3 1.35× ELDeR 50% 15.38 50.83 1801 1.75× 208.9 1.85× LLaMA2-13B Dense 0% 4.88 71.76 1066 1.00× 396.9 1.00× ELDeR 25% 5.82 65.11 1386 1.30× 298.4 1.33× ELDeR 50% 11.28 57.57 1823 1.71× 216.9 1.83×
-At a 50% pruning ratio on OPT-13B, ELDeR achieves a 75% increase in throughput and a 46% reduction in latency compared to the dense model.
+LLMs language processing involves two key phases with different bottlenecks: compute-bound prompt processing and memory-bound token generation.
+We measured the speedup for each stage individually.
+Table 2 presents the latency and throughput results for OPT-13B and LLaMA2-13B running on a single 80GB NVIDIA A100 GPU.
+Following prior method  Song et al.
+(2024), the token generation was tested by producing 128-token sentences with a batch size of 64, and prompt processing latency was assessed with a 2048-token input.
+Table 2: Throughput (tokens/s) and latency (ms) on OPT-13B and LLaMA2-13B. ‘PPL’ is the perplexity on Wikitext2. ‘PR’ is pruning ratio. ‘TI’ is throughput increase.
+Model Method PR PPL($\downarrow$) Avg_Acc(%) Tokens/s TI Latency($\downarrow$) Speedup OPT-13B Dense 0% 10.12 61.79 1029 1.00× 386.5 1.00× TRSP 25% 10.45 60.84 1348 1.31× 286.3 1.35× TRSP 50% 15.38 50.83 1801 1.75× 208.9 1.85× LLaMA2-13B Dense 0% 4.88 71.76 1066 1.00× 396.9 1.00× TRSP 25% 5.82 65.11 1386 1.30× 298.4 1.33× TRSP 50% 11.28 57.57 1823 1.71× 216.9 1.83×
+Pruning OPT-13B by 50% with TRSP yields a 75% increase in throughput and a 46% reduction in latency compared to the dense model.
 For LLaMA2-13B, it delivers a 71% improvement in throughput and a 45% decrease in latency.
-These results underscore ELDeR’s superior efficiency in accelerating model inference.
+These results underscore the end-to-end acceleration achieved by TRSP.
 4.4 Robustness to Different Pruning Ratios
-Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%.
-The perplexity of LLaMA2-7B under different methods are shown in Figure 4.
-ELDeR significantly outperforms other methods across various pruning ratios.
-At a pruning ratio of 20%, ELDeR performs similarly to the dense model.
-When the pruning ratio is up to 60%, SLEB collapses, while ELDeR maintains relatively low perplexity compared to other pruning methods.
-This demonstrates that ELDeR maintains robust performance even under extreme pruning ratios, enabling structured pruning to unlock significant potential for model acceleration.
-More detailed results are in Appendix E.
+Using the same settings as Section 4.2, we vary the pruning ratio from 20% to 60%.
+As shown in Figure 3, TRSP consistently achieves lower perplexity than other methods.
+At 20% pruning, it matches the dense model, and even at 60%, where SLEB fails, it maintains low perplexity.
+This demonstrates TRSP’s robustness and effectiveness in structured pruning for model acceleration.
+Additional results are provided in Appendix D.
+4.5 Dependency on Datasets
+Since TRSP relies on data-driven regularization, we investigate its dataset dependency.
+Keeping all other settings consistent with Section 4.2, we only change the source of the 128 samples: they are drawn respectively from Alpaca, WikiText-2, PTB, and C4.
+We evaluated perplexity of five methods on WikiText-2.
+As shown in Figure 4, TRSP consistently outperforms the other methods across datasets, demonstrating its robustness.
 Figure 3: Perplexity of LLaMA2-7B pruned by five methods under different pruning ratios on WikiText-2.
 Figure 4: Comparison of perplexity on different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
-4.5 Dependency on Datasets
-Since ELDeR relies on data-driven regularization, we investigate its dataset dependency.
-Keeping all other settings consistent with Section 4.2, we only change the source of the 128 samples: they are drawn respectively from Alpaca, WikiText-2, PTB, and C4.
-We evaluated perplexity of five methods on WikiText-2.
-As shown in Figure 4, ELDeR consistently outperforms the other methods across datasets, demonstrating its robustness.
-4.6 Low Overhead
-We evaluated the perplexity of baselines on LLaMA2-7B using varying amounts of data for RFT, keeping all other conditions consistent with Section 4.2 and only change the data size from 1,000 to 8,000.
-ELDeR is finetuing-free.
-From Figure 5, it can be observed that ELDeR outperforms the other methods with approximately 4.000 RFT data, significantly reducing data overhead.
-We speculate the reason is that ELDeR iteratively learns layer weights and then applies regularization, allowing it to more accurately identify layers to prune before pruning, and subsequently transfer the information from those layers to the remaining parts of the model through regularization.
-This process reduces information loss caused by pruning, thus preserving model performance and lowering the cost associated with RFT.
+4.6 Minimal Cost
+We only vary the retraining data size from 1,000 to 8,000 and use the same settings as in Section 4.2.
+We evaluated the perplexity of baselines on LLaMA2-7B.
+TRSP is retraining-free.
+From Figure 5, we observe that TRSP outperforms the other methods with 4,000 retraining data, significantly reducing the cost.
+We speculate that TRSP iteratively learns layer weights and then applies regularization, allowing it to identify layers to prune more accurately, and then transfer the knowledge from those layers to the remaining layers of the model through regularization.
+This process reduces knowledge loss, thus preserving model performance and lowering the retraining cost.
+Figure 5: The perplexity on WikiText-2 using TRSP and baselines under different amounts of retraining data.
 4.7 Ablation Study
-Effect of Learning Layer Weights Iteratively ELDeR learns layer weights in a greedy and iterative manner.
+Table 3: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘w/o W’ denotes learning layer weights in one-shot, ‘w/o R’ means no regularization. The pruning ratio is 25%.
+Model Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B TRSP 7.08 0.00 60.57 0.00 w/o W 9.26 +2.18 56.19 -4.38 w/o R 10.15 +3.07 54.36 -6.21 LLaMA2-13B TRSP 5.82 0.00 65.11 0.00 w/o W 8.35 +2.53 59.36 -5.75 w/o R 9.47 +3.65 56.25 -8.86
+Effect of Learning Layer Weights Iteratively TRSP learns layer weights in a greedy and iterative manner.
 As shown in Table 3 (Row 3 and Row 6), replacing iterative layer weight learning with a one-shot approach leads to increased model perplexity and decreased accuracy, highlighting the importance of learning layer weights iteratively, which will be discussed in detail in Section 4.8.
 Effect of Applying Regularization As shown in Table 3 (Row 4 and Row 7), removing the regularization process results in increased model perplexity and decreased accuracy, demonstrating the effectiveness of applying regularization, which will be discussed in detail in Section 4.9.
-Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda_{1}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$ and $\lambda_{2}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$.
+With all other settings the same as Section 4.2, we evaluate the model’s performance by varying $\lambda_{1}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$ and $\lambda_{2}\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$.
 We perform a grid search over $\lambda_{1}$ and $\lambda_{2}$, with details provided in Appendix C.
 The optimal combination yielding the lowest perplexity on LLaMA2-7B is $\lambda_{1}=5\times 10^{-3}$ and $\lambda_{2}=10^{-3}$.
-Figure 5: The comparison of perplexity on WikiText-2 between ELDeR and other baseline methods under different amounts of RFT data.
-Table 3: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘w/o W’ refers learning layer weights in one-shot, ‘w/o R’ refers no regularization. The pruning ratio is 25%.
-Model Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B ELDeR 7.08 0.00 60.57 0.00 w/o W 9.26 +2.18 56.19 -4.38 w/o R 10.15 +3.07 54.36 -6.21 LLaMA2-13B ELDeR 5.82 0.00 65.11 0.00 w/o W 8.35 +2.53 59.36 -5.75 w/o R 9.47 +3.65 56.25 -8.86
 4.8 Choice of Learning Layer Weights
 In this section, we explore the effectiveness of (1) iteratively learning layer weights using a greedy strategy compared to (2) acquiring all layer weights at one-shot.
-We set the pruning ratio to 25% and conduct experiments on LLaMA2-7B (32 layers), keeping all settings consistent with Section 4.2 except for the method used to determine layer weights.
+We set the pruning ratio to 25% and conduct experiments on LLaMA2-7B (32 layers), using the same settings as in Section 4.2.
 The perplexity and average accuracy of the pruned models are shown in Table 4.
@@
 Figure 6: Results of selecting the eight lowest-weight layers using iterative and one-shot layer weight learning.
-4.9 Impact of Regularization
+4.9 Impact of Stage Two Regularization
 We keep other settings consistent with Section 4.2.
-After learning the layer weights iteratively, we consider two scenarios: (1) without regularization and (2) with regularization, followed by pruning.
-The perplexity and average accuracy under (1) and (2) on LLaMA2-7B are shown in Table 5.
+After learning the layer weights iteratively, we consider two scenarios: (1) without regularization and (2) with regularization, then prune.
+The perplexity and average accuracy under (1) and (2) on LLaMA2-7B are in Table 5.
 The model exhibits lower perplexity and higher average accuracy with regularization.
@@
 A smaller loss function indicates better model performance indicating that the regularization process mitigates the impact of pruning on the overall model.
-According to previous work [59], the output and input of each layer in LLMs exhibit high similarity.
+According to previous work Liu et al.
+(2023), the output and input of each layer in LLMs exhibit high similarity.
 This inherent similarity enables us to apply regularization on the difference between the input and output of certain layers using only a small portion of data.
@@
 The computation is performed on layers [3, 6, 9, 11, 13, 20, 23, 29], which are selected using the iterative method described in Section 4.8, both before and after regularization.
-$\text{CosSim}(\mathbf{X_{\text{in}}^{i}},\mathbf{X_{\text{out}}^{i}})=\frac{1}{bn}\sum_{k=1}^{b}\sum_{j=1}^{n}\frac{\left\langle\mathbf{X_{\text{in}}^{i}}[k,j,:],\ \mathbf{X_{\text{out}}^{i}}[k,j,:]\right\rangle}{\left\|\mathbf{X_{\text{in}}^{i}}[k,j,:]\right\|_{2}\cdot\left\|\mathbf{X_{\text{out}}^{i}}[k,j,:]\right\|_{2}}$ (6)
 As shown in Figure 7, the input-output similarity of these layers is already high before regularization and increases even further after regularization.
-The increased similarity in the layers with regularization indicates that the input undergoes less change after passing through these layers, suggesting that less information is retained in them.
-In the Appendix F, we illustrate the changes in similarity between the input and output of the layers that were not regularized, before and after regularization.
-The results in Figure 8 of Appendix F show that the similarity in these layers decreases, meaning that, after regularization, the input undergoes greater changes when passing through these layers than before.
-This demonstrates that the regularization process weakens the influence of the regularized layers and enhances the influence of the unregularized parts, suggesting that regularization process may facilitate the transfer of information from the regularized layers to the rest of the model.
-Table 5: The performance differences between (1) and (2).
+The increased similarity in the layers with regularization indicates that the input undergoes less change after passing through these layers, suggesting that less knowledge is retained in them.
+In the Appendix E, we illustrate the changes in similarity between the input and output of the layers that were not regularized, before and after regularization.
+The results in Figure 8 of Appendix E show that the similarity in these layers decreases, meaning that, after regularization, the input undergoes greater changes when passing through these layers than before.
+This demonstrates that the regularization process weakens the influence of the regularized layers and enhances the influence of the unregularized parts, suggesting that regularization process may facilitate the transfer of knowledge from the regularized layers to the rest of the model.
+Table 5: The performance differences of (1) and (2).
 Cases PPL($\downarrow$) Avg_Acc(%) (1) 7.08 60.57 (2) 10.15 54.36
@@
 5 Conclusion
-In this paper, we propose a new pruning paradigm: first apply regularization, then prune.
-Unlike previous paradigm that first prune and then apply RFT, ELDeR transfers critical information from the pruned parameter space to the remaining parts during regularization, effectively mitigating the irreversible performance degradation caused by information loss.
-ELDeR demonstrates superior performance in generation and zero-shot tasks, significantly outperforming existing methods.
-For instance, ELDeR surpasses the powerful ShotGPT by 20% in perplexity on LLaMA2-7B.
-On OPT-13B, under 25% sparsity, the average accuracy drops by only 1%, while achieving 1.35× speedup compared to the dense model.
-Moreover, ELDeR is finetuning-free, substantially reducing computational costs and the reliance on RFT.
-The new paradigm may provide insights for structured pruning in LLMs.
+We propose a novel structured pruning method, TRSP.
+By performing two-stage regularization, TRSP retains more knowledge and better preserves model performance compared to direct parameter elimination.
+TRSP surpasses existing layer-wise pruning methods in generation and zero-shot tasks.
+For example, it reduces perplexity by 20% compared to ShotGPT on LLaMA2-7B, under 25% sparsity, the average accuracy decreases by just 1%, while delivering a 1.35× acceleration over the dense model.
+TRSP is retraining-free, significantly lowering computational overhead and dependence on retraining.
+The novel structured pruning method offers potential guidance for pruning strategies in LLMs.
+Limitations
+TRSP has primarily been evaluated on autoregressive language models and its applicability to other architectures or tasks remains unexplored.
+In future work, we plan to explore the application of TRSP to other model architectures, such as convolutional neural networks (CNNs).
 Appendix A Proof of Proposition 3.1
@@
 In our experiments, we employ FP16 precision for all evaluated models, including Phi-2, OPT-2.7B, LLaMA3-8B, OPT-13B, LLaMA2-7B, and LLaMA2-13B.
-For all RFT configurations, we set the LoRA rank $r$ to 32, the scaling factor $\alpha$ to 10, and the sequence length to 2048.
-All other hyperparameters follow the default settings provided in the Hugging Face PEFT package [60].
+For all retraining configurations, we set the LoRA rank $r$ to 32, the scaling factor $\alpha$ to 10, and the sequence length to 2048.
+All other hyperparameters follow the default settings provided in the Hugging Face PEFT package Mangrulkar et al.
+(2022).
 We set the batch size to 64.
@@
 Precision LoRA Rank Scaling Factor Max Sequence Length Batch Size Learning Rate Early Stop Threshold Min Delta FP16 32 10 2048 64 2e-5 5 1e-4
-To ensure a fair comparison between ELDeR and other methods, we maintain consistency in the data used across all approaches.
-Specifically, the data used by ELDeR for learning layer weights and performing regularization is identical to the calibration data used by the baseline methods.
-Furthermore, we ensure that the data employed during the RFT process is consistent across all baseline methods, thereby enabling a controlled and equitable evaluation framework.
-Following previous works [16], for the comparison unstructured pruning methods like Mangnitude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by ELDeR for learning weights and regularization.
+To ensure a fair comparison between TRSP and other methods, we maintain consistency in the data used across all approaches.
+Specifically, the data used by TRSP for learning layer weights and performing the second stage regularization is identical to the calibration data used by the baseline methods.
+Furthermore, we ensure that the data employed during the retraining process is consistent across all baseline methods.
+Following previous works Song et al.
+(2024), for the comparison unstructured pruning methods like Mangnitude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by TRSP for learning weights and regularization.
 Appendix C Optimal $\lambda_{1}$ and $\lambda_{2}$ for LLaMA2-7B
@@
 $\lambda_{1}$ / $\lambda_{2}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ $5\times 10^{-2}$ $10^{-1}$ $10^{-5}$ 12.28 11.91 8.94 10.48 13.79 14.73 15.87 $10^{-4}$ 11.45 10.36 8.26 9.82 11.87 12.56 13.62 $10^{-3}$ 10.82 9.14 7.92 8.53 9.75 10.82 12.98 $5\times 10^{-3}$ 10.09 8.35 7.08 7.46 8.52 10.17 12.25 $10^{-2}$ 11.26 9.97 8.13 8.92 10.46 12.88 13.43 $5\times 10^{-2}$ 12.89 11.25 10.38 11.76 12.85 14,75 15.37 $10^{-1}$ 14.55 13.58 12.25 14.41 15.72 16.91 17.93
-Appendix D Results of Comparison with Unstructured Pruning Methods
-We compare ELDeR at a 25% pruning ratio with unstructured pruning methods employing 2:4 sparsity (where two out of every four consecutive parameters are set to zero) and present the experimental results in Table 8.
-For unstructured pruning methods, only calibration data is required.
-To ensure consistency, all methods utilize a total of 128 samples from WikiText-2 training set, randomly selected from WikiText-2.
-The models used in these experiments consists of the OPT and LLaMA model families.
-The results are shown in Table 8.
-It can be seen that ELDeR with a pruning ratio of 25% significantly outperforms existing unstructured pruning methods with a sparsity of 2:4 on generative and zero-shot tasks, demonstrating the powerful performance of ELDeR.
-The reason ELDeR outperforms unstructured pruning in comparisons lies in iteratively learning layer weights and the regularization process before pruning.
-The regularization process transfers important information hidden in the parameter space which will be pruned to the remaining parts of the model, preventing irreversible information loss and performance degradation caused by direct pruning.
-Table 8: Performance comparison with unstructured pruning methods.
-‘PPL’ refers to the perplexity on WikiText-2.
-The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy.
-The best result is highlighted in bold, and the second-best in underlined.
-ELDeR-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while ELDeR-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
-Model Method Pruning Ratio PPL ($\downarrow$) PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 Magnitude 2:4 269 56.42 41.79 34.12 31.38 16.19 35.98 Wanda 2:4 16.98 63.78 56.39 40.29 41.37 23.28 45.02 SparseGPT 2:4 13.96 66.52 54.71 46.48 45.72 24.89 47.66 ELDeR-$\ell_{2}$ 25% 13.18 70.54 60.27 46.35 51.59 27.36 51.22 ELDeR-$\ell_{1}$ 25% 13.12 70.65 60.13 46.24 51.76 27.55 51.27 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 Magnitude 2:4 978 61.98 47.26 42.19 40.53 20.18 42.43 Wanda 2:4 14.08 70.53 59.14 53.83 50.86 27.31 52.33 SparseGPT 2:4 12.25 72.18 59.74 56.26 52.34 29.11 53.93 ELDeR-$\ell_{2}$ 25% 11.26 73.15 62.26 60.58 55.93 30.37 56.46 ELDeR-$\ell_{1}$ 25% 11.49 72.89 63.09 60.23 56.21 30.16 56.52 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 Magnitude 2:4 1.21e4 62.37 50.28 47.57 42.26 20.67 44.63 Wanda 2:4 13.72 71.67 60.65 56.81 52.64 28.51 54.06 SparseGPT 2:4 12.89 73.65 61.08 58.47 55.36 30.58 55.83 ELDeR-$\ell_{2}$ 25% 10.45 74.15 64.47 68.82 61.55 35.23 60.84 ELDeR-$\ell_{1}$ 25% 10.32 74.58 64.37 68.45 61.29 34.87 60.71 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 Magnitude 2:4 56.75 63.21 59.75 55.41 58.27 26.32 52.59 Wanda 2:4 11.87 69.75 61.56 54.24 56.38 30.84 54.55 SparseGPT 2:4 9.62 71.45 63.82 57.34 59.81 33.75 57.23 ELDeR-$\ell_{2}$ 25% 7.08 72.48 67.52 60.45 62.69 39.73 60.57 ELDeR-$\ell_{1}$ 25% 7.17 72.08 67.93 60.42 62.38 39.89 60.54 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 Magnitude 2:4 8.75 69.38 61.56 49.43 61.82 30.65 54.57 Wanda 2:4 8.49 72.85 66.28 62.73 63.09 36.52 60.29 SparseGPT 2:4 7.15 74.72 67.68 64.23 65.19 38.58 62.08 ELDeR-$\ell_{2}$ 25% 5.82 74.56 69.34 64.79 71.25 45.63 65.11 ELDeR-$\ell_{1}$ 25% 5.89 75.06 69.18 64.96 71.43 45.26 65.18
-Appendix E Performance of ELDeR under Different Pruning Ratios and Datasets
-E.1 The Perplexity of ELDeR under Different Pruning Ratios and Datasets
-To systematically evaluate the performance of ELDeR on zero-shot tasks across different pruning rates, we adopt the experimental setup outlined in Section 4.2, in which 128 samples are randomly sampled from the WikiText-2 training set to guide both the iterative learning of layer weights and the regularization process.
+Appendix D Performance of TRSP under Different Pruning Ratios and Datasets
+D.1 The Perplexity of TRSP under Different Pruning Ratios and Datasets
+To systematically evaluate the performance of TRSP on zero-shot tasks across different pruning rates, we adopt the experimental setup outlined in Section 4.2, in which 128 samples are randomly sampled from the WikiText-2 training set to guide both the iterative learning of layer weights and the regularization process.
 Subsequently, we evaluate multiple large language models (LLMs) by measuring changes in perplexity across various generative task datasets, including WikiText-2, Alpaca, PTB, and C4, under pruning rates of 10%, 20%, 30%, 40%, 50%, and 60%.
-The detailed results, presented in Table 9, indicate that ELDeR exhibits greater robustness as model scale increases, suggesting that the proposed method effectively mitigates performance degradation in larger architectures.
-This highlights the scalability of ELDeR and its potential to maintain model efficiency under varying levels of sparsity.
-Table 9: Perplexity comparison of ELDeR with different pruning ratios.
+The detailed results, presented in Table 8, indicate that TRSP exhibits greater robustness as model scale increases, suggesting that the proposed method effectively mitigates performance degradation in larger architectures.
+This highlights the scalability of TRSP and its potential to maintain model efficiency under varying levels of sparsity.
+Table 8: Perplexity comparison of TRSP with different pruning ratios.
 We set the pruning ratios to 10%, 20%, 30%, 40%, 50%, and 60%, and test the perplexity of the OPT and LLaMA2 models on the generation task datasets Alpaca, WikiText-2, PTB, and C4.
-For ELDeR, we use the $\ell_{2}$-norm.
+For TRSP, we use the $\ell_{2}$-norm.
 Model Pruning Ratio WikiText-2($\downarrow$) Alpaca($\downarrow$) PTB($\downarrow$) C4($\downarrow$) OPT-2.7B Dense 12.46 11.64 17.97 14.32 10% 12.78 11.89 18.65 15.02 20% 12.96 12.15 19.38 16.36 30% 15.52 13.47 24.28 20.75 40% 19.67 15.82 32.25 25.41 50% 25.62 21.89 47.67 33.73 60% 35.62 32.19 59.87 47.38 OPT-6.7B Dense 10.85 10.27 15.77 12.71 10% 11.03 10.95 16.78 13.46 20% 11.48 11.53 17.96 15.19 30% 12.87 12.75 20.86 17.52 40% 14.87 13.39 26.12 21.63 50% 21.43 16.51 35.36 28.74 60% 29.63 22.16 48.04 40.25 OPT-13B Dense 10.12 9.46 14.52 12.06 10% 10.28 9.89 15.12 12.73 20% 10.39 10.37 16.51 13.45 30% 11.38 11.12 19.42 16.14 40% 12.67 12.33 24.62 20.79 50% 15.38 14.76 30.88 27.65 60% 28.23 20.98 39.38 36.14 LLaMA2-7B Dense 5.47 5.25 7.92 7.26 10% 5.58 5.31 8.12 7.47 20% 6.13 5.87 8.78 7.92 30% 8.26 7.64 9.58 8.85 40% 10.28 9.79 12.87 11.42 50% 14.58 13.14 19.52 15.96 60% 25.18 21.46 30.14 28.32 LLaMA2-13B Dense 4.88 4.63 7.16 6.73 10% 4.99 4.91 7.48 7.93 20% 5.34 5.26 8.25 8.75 30% 6.87 6.35 8.97 9.68 40% 8.95 7.92 10.08 11.26 50% 11.23 9.73 12.63 14.73 60% 15.74 12.52 17.82 19.45
-E.2 The Accuracy of ELDeR under Different Pruning Ratios on Zero-shot Tasks
-To systematically evaluate the performance of ELDeR on zero-shot tasks across different pruning rates, we adopt the experimental setup outlined in Section 4.2, in which 128 samples are randomly sampled from the WikiText-2 training set to guide both the iterative learning of layer weights and the regularization process.
+D.2 The Accuracy of TRSP under Different Pruning Ratios on Zero-shot Tasks
+To systematically evaluate the performance of TRSP on zero-shot tasks across different pruning rates, we adopt the experimental setup outlined in Section 4.2, in which 128 samples are randomly sampled from the WikiText-2 training set to guide both the iterative learning of layer weights and the regularization process.
 We assess the accuracy of different model configurations at pruning rates of 10%, 20%, 30%, 40%, 50%, and 60% across a diverse set of benchmark datasets, including PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c.
 The results, summarized in Table 9, provide insights into the impact of sparsity on zero-shot generalization.
-Notably, the analysis reveals that ELDeR maintains competitive performance even at higher pruning rates, demonstrating its effectiveness in preserving reasoning and commonsense understanding across different tasks.
-Table 10: Accuracy comparison of ELDeR with different pruning ratios.
+Notably, the analysis reveals that TRSP maintains competitive performance even at higher pruning rates, demonstrating its effectiveness in preserving reasoning and commonsense understanding across different tasks.
+Table 9: Accuracy comparison of TRSP with different pruning ratios.
 We set the pruning ratios to 10%, 20%, 30%, 40%, 50%, and 60%, and test the accuracy of the OPT and LLaMA2 models on the zero-shot task datasets PIQA, WinoGrande, HellaSwag, ARC-e and ARC-c.
-For ELDeR, we use the $\ell_{2}$-norm.
+For TRSP, we use the $\ell_{2}$-norm.
 ‘Avg_Acc’ represents the average accuracy.
 Model Pruning Ratio PIQA(%) WinoGrande(%) HellaSwag(%) ARC-e(%) ARC-c(%) Avg_Acc(%) OPT-2.7B Dense 74.81 61.01 60.58 54.42 31.14 56.39 10% 72.35 60.75 51.53 52.73 29.15 53.30 20% 71.47 60.46 48.72 51.95 28.04 52.13 30% 66.73 58.53 44.36 50.02 26.58 49.24 40% 63.57 55.32 41.63 47.79 25.24 46.71 50% 58.48 52.29 40.52 46.26 21.32 43.77 60% 52.85 49.59 38.46 43.24 16.07 40.04 OPT-6.7B Dense 76.39 65.19 67.16 60.14 34.64 60.70 10% 75.42 63.38 65.16 57.23 32.89 58.82 20% 74.58 62.25 61.46 55.98 31.75 57.20 30% 71.88 60.56 59.15 53.62 28.54 54.75 40% 66.98 57.25 53.66 49.42 25.73 50.61 50% 62.78 54.56 45.63 46.38 21.75 46.22 60% 54.35 50.26 41.32 42.69 18.23 41.37 OPT-13B Dense 76.82 64.80 69.81 61.87 35.67 61.79 10% 75.49 64.67 69.26 61.73 35.54 61.34 20% 74.89 64.59 68.95 61.62 35.41 61.09 30% 71.46 62.67 66.53 59.39 33.61 58.73 40% 68.52 60.39 63.57 56.12 28.83 55.49 50% 63.35 57.62 57.43 51.65 24.09 50.83 60% 56.28 53.06 51.69 47.26 21.87 46.03 LLaMA2-7B Dense 79.11 69.06 75.99 74.58 46.25 69.00 10% 77.62 68.45 70.25 69.85 43.42 65.92 20% 75.38 67.96 65.26 65.83 41.57 63.20 30% 71.29 64.77 57.68 60.45 38.65 58.57 40% 67.61 60.38 54.12 57.53 35.45 55.02 50% 61.87 56.42 51.62 53.45 30.28 50.73 60% 55.39 51.54 48.73 47.42 27.08 46.03 LLaMA2-13B Dense 80.47 72.22 79.39 77.48 49.23 71.76 10% 78.45 71.87 74.32 76.26 48.39 69.86 20% 76.12 70.91 70.06 74.68 46.51 67.66 30% 72.54 68.15 63.51 70.88 43.21 63.66 40% 71.45 65.52 62.28 69.13 41.79 62.03 50% 67.36 58.52 60.65 63.45 37.88 57.57 60% 61.89 55.26 58.06 57.62 31.74 52.91
-Appendix F Similarity Changes in Unregularized Layers
+Appendix E Similarity Changes in Unregularized Layers
 We plot the input-output similarity of the unregularized layers in LLaMA2-7B, as described in Section 4.9, in Figure 8.
@@
 This suggests that more information is being captured in these layers compared to before, implying that the regularization process may cause information to shift from the regularized layers to the unregularized ones.
-Figure 8: A comparison between existing layer-wise structured pruning methods and ELDeR. Deeper blue layer represents greater performance impact, the taller cylinder represents larger data volume.
-Appendix G Future Work and Limitations
-ELDeR has primarily been evaluated on autoregressive language models and its applicability to other architectures or tasks remains unexplored.
-In future work, we plan to explore the application of ELDeR to other model architectures, such as convolutional neural networks (CNNs).
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: We have provided the main claims of our research in the abstract and introduction.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We have provided the limitation and future work discussion in Section G.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: Proposition 3.1 is theoretical. We provide its corresponding assumptions and proofs in Appendix A.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: Please refer to Section 4.1 in the paper.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: The dataset we used in this work is public available and we will release our reproducible code after the author notification.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Please refer to Section B in the paper.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [N/A]
-  34. Justification: The reason error bars are not included is due to the high computational cost required to compute them.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We have provided the experimental computing resources of our paper in Section 4.1.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44.
-Justification: We have carefully reviewed the NeurIPS Code of Ethics and ensured full compliance, including preserving anonymity.
-Our study uses publicly available models and datasets, with proper citations to the original sources.
-No aspects of our work violate the ethical guidelines set by NeurIPS.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: We have provided the broader impacts of our work in Section 5 and Section G.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: There is no such risks in our work.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We ensure all papers and code used or relevant to our paper are properly cited.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: We totally use public benchmarks.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [Yes]
-  79. Justification: We have described the usage of LLMs in our core method in Section 4.1.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Figure 8: A comparison between existing layer-wise structured pruning methods and TRSP. Deeper blue layer represents greater performance impact, the taller cylinder represents larger data volume.
+Appendix F Example Appendix
+$\begin{split}\text{CosSim}(\mathbf{X_{\text{in}}^{i}},\mathbf{X_{\text{out}}^{i}})=\quad\quad\quad\quad\quad\quad\quad\quad\\ \frac{1}{bn}\sum_{k=1}^{b}\sum_{j=1}^{n}\frac{\left\langle\mathbf{X_{\text{in}}^{i}}[k,j,:],\ \mathbf{X_{\text{out}}^{i}}[k,j,:]\right\rangle}{\left\|\mathbf{X_{\text{in}}^{i}}[k,j,:]\right\|_{2}\cdot\left\|\mathbf{X_{\text{out}}^{i}}[k,j,:]\right\|_{2}}\end{split}$ (9)